### v2

# Population Protocols Revisited: Parity and Beyond

Abstract
The population protocol model is used to study the power of pairwise interactions between simple, indistinguishable entities, referred to as agents. In this model, each agent is equipped with limited storage, represented by a single state drawn from a predefined state space. In the probabilistic variant of population protocols adopted here, a random scheduler selects an interacting pair of agents uniformly at random at each step of the protocol. In this variant, in addition to efficient state utilisation, one is also concerned with the stabilisation time, determined by the number of interactions required to reach the final configuration, divided by the population size $n$. We say that a protocol is time-efficient if it stabilises in polylogarithmic (in $n$) time, and space-efficient if it utilises polylogarithmic state space. We also say that a protocol is stable if it stabilises with the correct answer with probability 1, and silent if agents do not change their states after stabilisation.
For nearly two decades, population protocols have been extensively studied, yielding efficient solutions for central problems in distributed computing, including leader election, and majority computation, a predicate type in Presburger Arithmetic closely tied to population protocols [5, 7]. Surprisingly, no protocols have achieved both time- and space-efficiency for congruency predicates, including parity computation, which are complementary in this arithmetic framework. This gap highlights a significant challenge in the field [16, 19, 21, 24]. To address this gap, we explore the parity problem, where agents are tasked with computing the parity of the given sub-population size. Then we extend the solution for parity to compute congruences modulo an arbitrary $m$.
Previous research on efficient population protocols has focused on protocols that minimise both stabilisation time and state utilisation for specific problems. In contrast, this work slightly relaxes this expectation, permitting protocols to place less emphasis on full optimisation and more on universality, robustness, and probabilistic guarantees. This allows us to propose a novel computing paradigm that integrates population weights (or simply weights), a robust clocking mechanism, and efficient anomaly detection coupled with a switching mechanism (which ensures slow but always correct solutions). This paradigm facilitates universal design of efficient multistage stable population protocols. Specifically, the first efficient parity and congruence protocols introduced here use both $O(\log^{3}n)$ states and achieve silent stabilisation in $O(\log^{3}n)$ time. We conclude by discussing the impact of implicit conversion between unary and binary representations enabled by the weight system, with applications to other problems, including the computation and representation of (sub-)population sizes, a problem posed in [21].
All algorithmic solutions presented in this paper are stable, silent, and guarantee stabilisation time with high probability (whp) defined as $1-n^{-\eta},$ for a constant $\eta>0.$
1 Introduction
The model of population protocols originates from the seminal paper [5], providing tools suitable for the formal analysis of pairwise interactions between simple, indistinguishable entities referred to as agents. These agents are equipped with limited storage, communication, and computation capabilities. When two agents engage in a direct interaction, their states change according to the predefined transition function, which forms an integral part of the population protocol. The weakest assumptions in population protocols concern the fixed (constant-size) operational state space of agents, with the population size $n$, whether exact or approximate, not hard-coded into either the states or the transition function. It is assumed that a protocol starts in the predefined initial configuration of agents’ states representing the input, and it stabilises in one of the final configurations of states representing the solution to the considered problem. In the standard probabilistic variant of population protocols adopted here, in each step of a protocol, the random scheduler selects an ordered pair of agents: the initiator and the responder, which are drawn from the whole population uniformly at random. The lack of symmetry in this pair is a powerful source of randomness utilised by population protocols. In the probabilistic variant, in addition to efficient state utilisation, one is also interested in the time complexity, where the sequential time refers to the number of interactions leading to the stabilisation of a protocol in a final configuration. More recently, the focus has shifted to the parallel time, or simply the time, defined as the sequential time divided by the size $n$ of the whole population. The (parallel) time reflects on the parallelism of simultaneous independent interactions of agents utilised in efficient population protocols that stabilise in time $O(\mbox{poly}\log n)$. All algorithmic solutions presented in this paper are stable (always correct), silent (agents do not change their states after stabilisation), and guarantee stabilisation time with high probability (whp) defined as $1-n^{-\eta},$ for a constant $\eta>0.$
In this paper, the rules of a transition function are denoted as $A+B\rightarrow A^{\prime}+B^{\prime}$, meaning that if there is an interaction initiated by any agent in state $A$, with any agent in state $B$, then the former changes its state to $A^{\prime}$, while the latter changes its state to $B^{\prime}$. Some of the protocols presented in this work operate on compound states, which means that an agent state is a concatenation of some sets of substates $S_{1},\ldots,S_{k}$, and defining these sets defines the full state space as $S=S_{1}\times\ldots\times S_{k}$. Encoding a state of an agent can be seen as labeling it by a number represented in a binary form. Encoding compound states can be achieved by utilising specified positions in binary representation for each substate set.
1.1 Related Work and Results
Population protocol studies primarily follow two main research avenues. The first focuses on understanding the behavior of large populations driven by simple interaction rules. The second treats population protocols as computational systems, investigating their computational power based on factors such as the type of scheduler and the memory available to individual agents. This study refers to the latter, focusing on designing efficient population protocols with robust probabilistic guarantees.
The original population protocol model [5] assumes that the number of states and the transition function are constant, i.e., independent of the population size $n$. However, for fundamental problems such as leader election [23], majority computation [1], and the evaluation of various functions and predicates [9], it has been demonstrated that stable, fixed-state protocols operating in sublinear time are infeasible. This limitation has motivated the study of population protocols in which the number of states is permitted to grow with $n$, enabling such problems to be solved efficiently, i.e., in polylogarithmic time.
Prior work on efficient population protocols has emphasized minimizing both stabilisation time and state usage for specific problems. In contrast, this study adopts a broader approach, prioritizing universality, robustness, and probabilistic guarantees over full optimisation. We introduce a novel computing paradigm that incorporates population weights (or simply weights), a robust clocking mechanism, and efficient anomaly detection paired with a switching mechanism to ensure slow but consistently correct solutions. This new paradigm supports the universal design of efficient multistage protocols, which are stable (always correct), silent (agents don’t alter their states after stabilisation), and ensure stabilisation in time O(polylog n)O($\text${polylog }n) with high probability.
In the remainder of this section, we present an overview of both established and emerging concepts that underpin the new paradigm and its associated algorithmic solutions. We also situate our contributions within the broader research context.
Epidemic Process
A (one-way) epidemic process [6] is a simple yet effective protocol in which each agent can be in one of two possible states, denoted as 0 (susceptible) or $1$ (infected), with the transition function consisting of a single significant rule $(x)+(y)\rightarrow(x)+(\max\{x,y\})$.
Lemma 1 (Lemma 2 in [6]).
The time required for all agents to become infected, starting from a configuration with a single infected agent, is $\Theta(\log n)$ whp.
The epidemic process establishes a fundamental $\Theta(\log n)$-time communication bound in population protocols. It lies at the core of numerous algorithmic tools and solutions in population protocols, including phase clocks [2, 6, 8, 22, 26] and, in turn, fast leader election [3, 15, 2, 14, 8, 25, 13], majority computation [4, 1, 2, 15, 29, 30, 11, 10, 22], and others. A recent paper [27] presents a generalisation of the epidemic process to the $k$-contact epidemic model, in which an agent becomes infected only after interacting with $k$ distinct infected agents.
Efficient Leader Election
A central challenge in distributed computing is symmetry breaking, typically achieved by electing one agent as the leader. In the context of population protocols, the results in [17, 20] laid the foundation for proving that leader election cannot be solved in sublinear time using a fixed number of states [23]. Alistarh and Gelashvili [3] then introduced an upper bound with a protocol stabilising in $O(\log^{3}n)$ time using $O(\log^{3}n)$ states per agent. Later, Alistarh et al. [1] explored general trade-offs between state and time complexity, establishing a separation between slowly stabilising protocols using $o(\log\log n)$ states and rapidly stabilising ones requiring $O(\log n)$ states. Subsequent advances improved these bounds: Bilke et al. [15] achieved $O(\log^{2}n)$ time whp with $O(\log^{2}n)$ states; Alistarh et al. [2] and Berenbrink et al. [14] reduced states to $O(\log n)$ while preserving $O(\log^{2}n)$ time whp using synthetic coins; Gąsieniec and Stachowiak [8] further lowered state usage to $O(\log\log n)$ with the same time bound whp; Gąsieniec et al. [25] traded high probability for expected $O(\log n\log\log n)$ time; and finally, Berenbrink et al. [13] achieved the optimal expected $O(\log n)$ time.
In this paper, a direct use of leader election via pairwise duels appears in the slow parity protocol, see Section 2.1. In fast protocols, the elected leader initiates weight creation and coordinates rounds and process synchronisation. Our leader election mechanism builds on coin-flip reduction across successive rounds [6, 8, 13], with round numbering ensuring that at least one leader survives. In the rare event of multiple surviving leaders, anomaly detection resolves the conflict in expected $O(n)$ time, see Section 3.3.
Majority Computation
Let $A$ and $B$ be subpopulations of $n$ agents. The majority problem is to determine whether $|A|>|B|$, $|A|<|B|$, or $|A|=|B|$. For efficient majority computation, Angluin et al. [6] introduced an approximate majority protocol: if the initial bias $|A|-|B|=\omega(\sqrt{n}\log n)$, it converges to the majority opinion in $O(\log n)$ time whp. Condon et al. [18] improved the required bias to $\Omega(\sqrt{n\log n})$. For the exact majority, Alistarh et al. [4] gave the first stable polylogarithmic-time protocol using $\Omega(n)$ states. Subsequent work [1, 2, 15, 29, 30, 11, 10] improved the space-time trade-off, culminating in an optimal $O(\log n)$-time protocol with $O(\log n)$ states [22]. This matches the $\Omega(\log n)$ state lower bound of Alistarh et al. [2] for stable sublinear-time protocols under monotonicity and output dominance, both satisfied by [22]. Non-stable protocols (allowing positive error probability) achieve faster convergence whp with less memory. Berenbrink et al. [11] gave a protocol with $O(\log\log n)$ states and $O(\log^{2}n)$ time. Also, Kosowski and Uznański [28] achieved $O(1)$ states and $O(\mathrm{polylog}\,n)$ time whp.
In this paper, we use exact majority computation as a conceptual balance scale, see Figure 2, enabling comparison of sub-population cardinalities through appropriately chosen weight combinations. Like other modern majority protocols, ours also uses cancelling (opposite biased agents $\to$ unbiased) and splitting (biased $\to$ two biased), doubling the bias per phase until $n$. As indicated earlier, our protocols prioritise universality, robustness, and probabilistic guarantees over full optimisation. For details, see Section 3.4.
Figure 1: Balance scale. [^2] This is an AI-generated picture.
Clocks and Synchronisation
Efficient population protocols typically use phase clocks to synchronise computation. Most such clocks are designed to count $\Theta(\log n)$ parallel time, which, assisted by epidemic propagation, is sufficient to fully synchronise all agents in the population. Leader- or junta-based phase clocks operate with a fixed number of states [6, 8], and when nested, they can count any polylogarithmic duration. An alternative class consists of leaderless clocks that utilise $\Theta(\log n)$ states [2, 22], enabling approximate parallel-time counting either as fixed-resolution clocks [22] or as oscillators [2]. A more recent clock type, introduced in [26], supports counting $\Theta(n\log n)$ parallel time using a fixed number of states, leveraging either leaders or network connections in the network constructor model.
In this paper, synchronisation is achieved via round counting by each agent, reinforced by leader-attributed fine-grained interaction counting and epidemic waves. This precise time-stamping system enables protocols with efficient anomaly detection through execution-time inconsistencies, essential for the MC+ protocols.
MC+ Protocols
Our new computing paradigm relies on Monte Carlo protocols enhanced with anomaly detection, termed MC+ protocols. These either silently stabilise to the correct answer or detect efficiently an anomaly, triggering a slow (polynomial-time) but always correct solution. We distinguish between clock anomalies, caused by desynchronisation, and protocol anomalies, which arise from interactions between agents in conflicting states. We show that MC+ protocols are suitable building blocks for constructing more complex solutions. For detail see Section 3.
Further Context
In the population protocol model, integer values are represented in unary format, where each value corresponds to the cardinality of a subpopulation of agents in a designated state. This representation makes addition straightforward, i.e., fast and stable. While it is well understood that fixed-state Monte Carlo population protocols can simulate a limited variant of register machines (recognising problems in LOGSPACE) [6], the efficient and stable computation of basic operations such as subtraction (multiplication and division) remains significantly more challenging. The primary obstacle in subtraction lies in the need for effective pairing of agents from two subpopulations, which may require linear time when their cardinalities are comparable or equal. A similar challenge arises in majority protocols, where it is addressed by amplifying the difference (bias) between the cardinalities to a magnitude that population protocols can process efficiently [6]. In the same paper, the authors propose a method for efficient Monte Carlo type subtraction by iteratively generating powers of two, combined with simple addition and more sophisticated majority protocols. This approach serves as a prototype for our system of weights, which introduces an implicit binary representation of integer values, supported by corresponding operations.
Our Contribution
While this paper primarily focuses on the parity problem, our new computing paradigm also enables efficient computation of congruences, by providing a suitable framework which facilitates universal design of efficient multistage stable population protocols (Section 3.2). Specifically, the first efficient parity (Theorem 1) and congruence (Theorem 9) protocols are proved correct, both using $O(\log^{3}n)$ states and achieving stabilisation in $O(\log^{3}n)$ time (Section 4). Both protocols are stable (always correct), silent (agents do not change their states after stabilisation), and guarantee stabilisation time whp. We conclude (Section 5) by discussing the impact of implicit conversion between unary and binary representations enabled by the weight system, with applications to other problems, including computation and representation of (sub-)population sizes.
To streamline the presentation, we assume that the subpopulations of agents representing input, temporary, and output values are each at most $\frac{n}{6}$. However, our solutions remain correct and efficient even for larger fractions of $n$. We also assume that the upper estimate $LOG$ on $\log n$ is provided with constant-factor accuracy and incorporated into the transition function. Finally, we note that a recent study [19] constructs succinctly encoded slow (polynomial time) population protocols for Presburger arithmetic, including congruences modulo any integer $m$. However, since [19] does not address the silence property, for completeness we include here a natural extension of our parity protocol that achieves silence.
2 The Parity Problem
Recall that in the parity problem, agents are tasked with computing the parity of a given sub-population size. In this work, we present a stable population protocol that solves the parity computation problem in polylogarithmic time. A common strategy for designing such protocols involves running a fast Monte Carlo algorithm that solves the problem with high probability, complemented by a slower, stable backup when needed. This approach has proven effective for problems such as majority computation and leader election, as discussed in the introduction. We adopt the same methodology in our construction. Notably, our protocol ensures both stabilisation and silence.
We begin (Section 2.1) by presenting a simple protocol that computes the parity of a given subpopulation $X$ of agents, using a fixed number of states. It always stabilises in time $O(n\log n)$, both in expectation and with high probability. To our knowledge, it represents the current state-of-the-art for stable parity computation and serves as the backup in our overall protocol design.
2.1 Slow Stable Parity Protocol
The slow protocol builds on the leader election protocol of [5]. It uses four states: $L_{0},L_{1},F_{0},F_{1}$}, with $L_{x}$ denoting leaders and $F_{x}$ followers. Initially, all agents in subpopulation $X$ start in $L_{1}$; all others start in $F_{0}$. On the interaction of two leaders, the rule $L_{x}+L_{y}\rightarrow L_{(x+y)\bmod 2}+F_{0}$ applies, where one remains a leader (with updated parity), and the other becomes a follower. Eventually, a single leader remains, encoding the parity of $|X|$ in its state index. This is then propagated via rule $L_{x}+F_{y}\rightarrow L_{x}+F_{x}$, so all followers adopt the leader’s index.
Lemma 2.
The slow protocol computes parity and stabilises silently in expected time $O(n\log n)$.
Proof.
The leader election protocol, obtained by removing indices from the parity protocol, is stable and stabilises in expected $O(n)$ time [5]. The parity of $|X|$ is encoded in the final leader’s state index: $x\equiv|X|\pmod{2}$. By a Coupon Collector argument, propagating this information to all agents takes expected $O(n\log n)$ time. Thus, the full protocol runs in expected $O(n\log n)$ time. ∎
2.2 Fast Parity Protocol
In this section, we provide an outline of the first fast, stable, and silent parity protocol (Algorithm 1).
1: $X$, a set of agents
2:true if $|X|$ is even, otherwise false
3:function Parity($X$)
4:  LeaderElection
5:  $\textsc{WeightCreation}(X)$
6:  $Y\leftarrow\emptyset$
7:  for $\ell\leftarrow LOG$ downto 11 do
8:   Add weight $2^{\ell}$ to $Y$
9:   Compare $X$ and $Y$ using the Majority protocol
10:   if $|X|<|Y|$ then Remove weight $2^{\ell}$ from $Y$
11:  if $|X|=|Y|$ then return true else return false
Algorithm 1 Parity Protocol
Balance scale
The balance scale idea stems from recomputation of powers of 2 (due to limited space), as proposed in [6]. Equipped with polylogarithmic memory we can afford initial construction of subpopulations $W_{i}$ of cardinality $2^{i},$ for all $i=0,1,\dots,\lfloor\log\frac{n}{3}\rfloor$. These subpopulations serve as weights in balance scales (Figure 2) implemented via the fast majority protocol. Using subpopulation $W_{i}$, we can determine whether $|X|-|Y|\geq 2^{i}$ by computing exact majority (with tie detection) between $|X|$ and $|Y|+|W_{i}|$. This balancing mechanism enables efficient computation of parity and size of any subpopulation.
Implementing this idea as a Monte Carlo protocol is rather straightforward; the challenge lies in designing mechanisms to ensure stable and silent convergence to the correct solution. This paper implements the weight-based parity computation via a Monte Carlo protocol using LeaderElection, WeightCreation, and Majority subprotocols. The LeaderElection protocol is efficient and is required for the correct execution of the other two subprotocols. The WeightCreation protocol creates weights $W_{i}$ and the Majority decides whether $|X|>|Y|,|X|<|Y|$ or $|X|=|Y|$, for any subpopulations $X$ and $Y$. Our efficient protocol either stabilises silently on the correct parity or detects an anomaly under unusual behavior. In the latter case, it invalidates its result and triggers the slow parity protocol. In Section 3, we develop the tools required to prove the following theorem.
Theorem 1.
There exists a population protocol which solves the parity problem utilising $O(\log^{3}n)$ states and stabilises silently in expected $O(\log^{3}n)$ time.
The proof of Theorem 1 is provided in Section 3.6, near the end of the main part of this paper.
3 Efficient Protocols with Anomaly Detection
In this section, we develop methods for constructing efficient population protocols that silently stabilise with certainty. Specifically, assuming the existence of two protocols: a slow, state-efficient protocol that stabilises silently in polynomial time with probability 1, and a fast Monte Carlo type protocol possibly composed of subprotocols, we show how to integrate these two to obtain a protocol that is both efficient in time and guaranteed to stabilise silently.
Monte Carlo protocols typically aim to compute the desired solution with high probability, but the correctness of the solution remains uncertain after execution. Existing mechanisms for detecting fast protocol failures and transitioning to a consistently stabilising protocol have been developed on an ad hoc basis. In contrast, we propose a novel universal anomaly-detection mechanism that effectively integrates multiple Monte Carlo subprotocols with a reliable, though slower, always-correct solution. Specifically, this mechanism identifies efficiently scenarios necessitating a shift to a slower, always-stabilising protocol and facilitates the creation of complex population protocols by sequentially combining simpler ones. Additionally, the constructed protocols are silent.
Our efficient (polylogarithmic time and space) protocols converge to the correct solution with high probability or signal potential errors via an efficient anomaly-detection mechanism. Such signal is treated as a failure of the fast protocol, even if it does not always indicate an actual error. Our universal strategy is to first run a fast Monte Carlo protocol that either stabilises with high probability or, upon detecting an anomaly, switches to a slower, always-stabilising fallback.
3.1 Clock Protocol
Central to our design is the clock protocol, which acts as a clocking device and a command counter, tracking algorithmic progress and synchronizing all agents to the same point in time. In rare cases of desynchronisation, it signals an anomaly. Each main (e.g., parity) protocol is built by sequentially composing subprotocols, each spanning one or more rounds. The clock counts down $R=\text{poly}\log(n)$ rounds of the main protocol, each taking $O(\log n)$ time. To measure a round’s duration, a leader-maintained interaction counter tracks $c\log n$ interactions, triggering the transition to the next round.
Components of the clock protocol
  • Each agent maintains a round number $r$.
  • Leaders maintain an auxiliary counter $s$ to count interactions within a round. After $c\cdot LOG$ interactions, they increment $r$ by 1, advancing to the next round until $r=R$.
  • If two agents with round numbers $r$ and r1r-1 interact, the agent with r1r-1 adopts $r$.
  • If two agents’ round numbers differ by more than 1, a clock anomaly is signaled.
Lemma 3.
For any constant $C>0$, there exists a constant $c>0$ such that counting $c\log n$ interactions takes at least $C\log n$ time whp. Moreover, there exists a constant $C_{0}>0$ such that, within time $C_{0}\log n$ after the first leader enters round $r$, all other agents advance to round $r$ whp.
Proof.
Each agent experiences on average $2C\log n$ interactions in time $C\log n$. By the Chernoff bound, this number does not exceed $c\log n$ whp, for sufficiently large $c$. The epidemic time to propagate round $r$ is at most $C_{0}\log n$ whp, for some $C_{0}>0$, ensuring all agents adopt round $r$ within this time. ∎
In the clock protocol, we choose $c>0$ such that counting $c\log n$ interactions takes time $C\log n$ whp, ensuring that the interval $(C-C_{0})\log n$ suffices for all round-$r$ computations whp. By Lemma 3, all agents occupy round $r$ simultaneously for at least $(C-C_{0})\log n$ time whp. Also by Lemma 3, round numbers differ by at most 1 across any two agents throughout all $R$ rounds whp. Thus, no clock anomaly is detected whp.
Lemma 4.
The expected time for the clock to terminate after $R$ rounds, either by completion or anomaly detection, is $O(R\log n)$.
Proof.
Let us consider a modified clock protocol in which anomaly detection is replaced by a new rule. This rule specifies that when two agents interact in rounds $r$ and $r^{\prime}$, with $r^{\prime}<r$, the agent in round $r^{\prime}$ adopts round $r$. Observe that the anomaly-detection clock protocol terminates in time no worse than the modified one. The two protocols either complete their counting during the same time, or, if an anomaly is detected, the original clock finishes earlier than the modified one. To bound the time for the modified clock protocol to terminate after $R$ rounds, consider the computation of an agent that remains a leader throughout. For the modified clock to finish, firstly, this agent must count $R$ rounds. Counting each of them takes no more than $c\log n$ interactions of the leader. Hence, counting $R$ rounds takes the leader expected time $O(R\log n)$. Secondly, all the remaining agents must also advance to round $R$, which occurs in time no worse than the time of the epidemic process by which the leader, already in round $R$, spreads round $R$ to the population. This takes the expected time $O(\log n)$. Therefore, the modified clock protocol terminates in expected time $O(R\log n)$. As noted earlier, this also implies that the anomaly-detection clock protocol terminates in expected time $O(R\log n)$. ∎
3.2 MC+ Protocols
This part demonstrates how to effectively integrate multiple population protocols, necessitating a more sophisticated state space structure. In particular, agent states are recorded as tuples $(A_{0})(A_{1})(A_{2})(A_{3})$. Element $A_{0}$ contains the input of the protocol, i.e., set $X$. Element $A_{1}$ contains global information computed by the protocol, i.e., whether the agent is a leader or belongs to one of the weights and whether this weight is currently placed on the scale pan. Element $A_{1}$ will also contain the output. We call this element the global state of an agent. Element $A_{2}$ specifies the round number $r$ and contains the leader’s interaction counter in a given round. Element $A_{3}$ contains local information specific to the subprotocol executed in round $r$. We call this part the local state of an agent. The transition $(A_{0})(A_{1})(A_{2})(A_{3})+(B_{0})(B_{1})(B_{2})(B_{3})\rightarrow(A_{0}^{\prime})(A_{1}^{\prime})(A_{2}^{\prime})(A_{3}^{\prime})+(B_{0}^{\prime})(B_{1}^{\prime})(B_{2}^{\prime})(B_{3}^{\prime})$ is implemented as follows. First, we use the clock transition rule to calculate the round number in states $A_{2}^{\prime}$ and $B_{2}^{\prime}$, optionally incrementing the interaction counter(s) in the round. Simultaneously, we calculate other states $A_{2},B_{2}$ according to the transition rules for the subprotocol. If necessary, we also update the information whether the agent is still the leader or about the weights or output in $A_{1}^{\prime},B_{1}^{\prime}$. Elements $A_{0},B_{0}$ remain unchanged.
In this work, we focus on fast population protocols, specifically Monte Carlo protocols with anomaly detection (MC+ protocols). These protocols either silently stabilise on the correct computational output whp or trigger an anomaly detection signal. To maintain order in protocol analysis, we distinguish two types of anomalies:
  • clock anomaly, that arises when agents whose rounds differ by more than 1 interact,
  • protocol anomaly, which involves the detection of irregularities specific to the protocol itself.
To capture the desired properties, we introduce the definition of MC+ protocols.
MC+ protocol
  • Is active for a specified number of rounds, meaning its operations are limited to $R$ consecutive rounds. Note that if there are no anomalies, this means reaching a silent configuration.
  • There is a negligible probability that a protocol anomaly signal will be triggered during the protocol’s execution or within an expected time of $O(n^{a})$ following its completion. Such an anomaly signal would stem from a state inconsistency, potentially indicating erroneous behavior in the protocol.
  • The absence of such a signal, together with the lack of clock anomaly, indicates a successful execution of the task.
Theorem 2.
If a fast MC+ protocol and a polynomial-time stable protocol exist for the same problem, they can be combined to yield a protocol that is both fast (in expected running time) and stable (with high probability). Moreover, if the polynomial-time stable protocol is silent, the resulting combined protocol is also silent.
Proof.
Assume that the fast protocol completes its computations in $R$ rounds. If not successful, it reports a protocol anomaly in at most $n^{a}$ average time with negligible probability less than $n^{-a}$, for suitably chosen protocol parameters and $a>0$. In this time, by Lemma 3 the clock reports an anomaly with negligible probability less than $n^{-a}$. Assume also that the slow algorithm runs in expected time less than $n^{a}$. The expected execution time of the combined algorithms consists of the following components:
  • the expected clock completion time which by Lemma 4 is $O(R\log n)$,
  • the execution time of the slow algorithm times the probability of clock anomaly: $n^{a}\cdot n^{-a}=1$,
  • the average time needed to handle protocol anomaly times its probability: $(n^{a}+n^{a})\cdot n^{-a}=2$.
Therefore, the expected execution time of the combination of fast and slow protocol is at most $O(R\log n)+1+2=O(R\log n).$ Moreover, the combination of the protocols is stable and silent.
Theorem 3.
A sequence of a polylogarithmic number of MC+ protocols is an MC+ protocol.
Proof.
Assume that there is a sequence of $k$ MC+ protocols running in $R_{1},R_{2},\ldots,R_{k}$ rounds. There exists $a>0,$ s.t., the probability of detecting a protocol anomaly is at most $n^{-a}$ for each protocol, and the time needed to detect an anomaly is, on average, less than $n^{a}$. So, the probability of detecting a protocol anomaly in a series of these protocols is $kn^{-a}$, which is negligible. The total number of rounds is $R=\sum_{i}R_{i}$, which is polylogarithmic. The average time needed to detect a protocol anomaly after completion of the protocol is at most $n^{a}$. ∎
3.3 Leader Election
The leader election protocol proposed in this work builds upon a scheme that has been employed in several prior studies [6, 8, 13]. In this approach, the leaders perform coin flips in successive rounds. Whether an agent draws heads or tails depends on the random scheduler’s designation of that agent as either the initiator or the responder in an interaction. If, in a given round, one leader flips tails while another flips heads, the leader obtaining tails becomes a follower. Unlike previous protocols, our method associates the coin-flip outcome (heads or tails) with the round number, thereby ensuring that at least one leader always persists. Apart from clock anomalies that can be detected during the execution of the protocol, the only other potential failure arises when multiple leaders remain. This is identified as a protocol anomaly when two such leaders interact after the leader election has concluded. This detection occurs in the expected time $O(n)$.
1: $X$, a set of agents
2:A set of agents with a single leader
3:function LeaderElection($X$)
4:  All agents adopt state $L$, indicating that each is a leader (candidate)
5:  for round $r\leftarrow 0$ to $d\cdot LOG$ do
[…]
7:   if an agent flipped H as a leader or received information that H was flipped in $r$ then
8:      It spreads by epidemic information that H was flipped in $r$
9:   if a leader flipped T and received information that H was flipped in $r$ then
10:      This leader changes type to a follower
11:  if after protocol end there is an interaction between two leaders then signal Anomaly
Algorithm 2 LeaderElection Protocol
Theorem 4.
There exists a constant $d>0$ such that LeaderElection (Algorithm 2) is an MC+ protocol, which either elects a single leader whp or reports an anomaly. Moreover, protocol anomaly is detected in the expected time $O(n)$ following the completion of the protocol.
Proof.
Let $A_{r}$ denote an event in round $r$, defined as follows. If round $r$ has at least two leaders, $A_{r}$ is the event that at least half, but not all, of the leaders flip tails. In this case, the probability of $A_{r}$ is at least $3/8$ (with the minimum occurring when there are three leaders). When event $A_{r}$ occurs, since $c$ is chosen to be sufficiently large, the epidemic in round $r$ succeeds whp, ensuring that at least half of the leaders are eliminated. If round $r$ has exactly one leader, $A_{r}$ is defined as the event that an independent draw with success probability $3/8$ succeeds.
A sufficient condition for electing a single leader is that event $A_{r}$ occurs in at least $\log n$ rounds. Over $d\cdot LOG$ rounds, the expected number of occurrences of $A_{r}$ is at least $\frac{3}{8}d\log n$. By the Chernoff bound, for sufficiently large $d$, event $A_{r}$ occurs at least $\log n$ times whp. Hence, a single leader is elected whp. If this is not the case, an anomaly is signaled in an average time $O(n)$ required for two remaining leaders to interact. For pseudocode of Algorithm see page 2. ∎
3.4 Exact Majority
We now construct and analyse our majority computation protocol. This protocol operates over $LOG+4$ rounds and achieves a time complexity of $O(\log^{2}n)$. Throughout this construction, we assume that a unique leader has already been selected. The protocol compares two subpopulations, $X$ and $Y$, determining which is larger or whether they are of equal size. Notably, $X$ and $Y$ need not be disjoint. At the start of the first round, agents are assigned states as follows: elements exclusive to $X$ are assigned state $(1)$; elements exclusive to $Y$ state (1)(-1); elements in both $X$ and $Y$ or in neither state $(0)$. During each round $r$, an averaging operation is performed on pairs of agents in round $r$: $(a)+(b)\rightarrow\left(\lfloor\frac{a+b}{2}\rfloor\right)+\left(\lceil\frac{a+b}{2}\rceil\right)$. When transitioning to the next round, a doubling operation is applied: $(a)\rightarrow(2a)$.
If, when moving to round $r$, we try to perform doubling operation and it yields state (a)=(2|2)(a)=(-2|2) (that is, (a)=(2)(a)=(-2) or $(a)=(2)$), the following transition is performed: $(2)\rightarrow(x_{win})$ or $(-2)\rightarrow(y_{win})$. Once agents in state $(x_{win})$ or $(y_{win})$ appear, the corresponding state is propagated by the epidemic process: $(x_{win})+(0|1|2)\rightarrow(x_{win})+(x_{win})$ or $(y_{win})+(-2|-1|0)\rightarrow(y_{win})+(y_{win})$. Note that e.g. states (1|2)(-1|-2) are not susceptible to infection by state $(x_{win})$. In this scenario, when moving on to round $r+1$, all agents are in state $(x_{win})$ or $(y_{win})$ whp. The leader is the first to advance to round $r+1$ by and makes whp the transition: $(x_{win})\rightarrow(X_{win})$ or $(y_{win})\rightarrow(Y_{win})$. Subsequently, all other agents moving to round $r+1$ perform exactly the same transition as the leader whp. The state $(X_{win})$ or $(Y_{win})$ remains unchanged until the final round of the protocol, unless an anomaly is signaled. There are two unlikely cases when an anomaly is detected. The first such case is when one of the agents has a state different from the leader’s state $(x_{win})$ or $(y_{win})$ when moving to round $r+1$. The second case is that the leader had a state other than $(x_{win})$ or $(y_{win})$ when entering round $r+1$, while another agent entering that round has one of these states.
If $X$ and $Y$ have the same cardinality, the protocol avoids states $(x_{win})$ and $(y_{win})$ whp. If, at the end of round $LOG+3$, the leader is in state (2|1|0|1|2)(-2|-1|0|1|2), it adopts state $(tie)$ in the next round. This state $(tie)$ is then propagated by the epidemic process to all agents entering round $LOG+4$ in states different to $(x_{win})$ and $(y_{win})$. However, if an agent enters round $LOG+4$ in state $(x_{win})$ or $(y_{win})$, a protocol anomaly is detected.
In the proofs concerning our majority protocol, we make use of a Theorem from [12]. Below, we provide an equivalent formulation of this result, adapted to the model employed in our work.
Theorem 5 (Theorem 1 in [12]).
Suppose we are given an initial configuration in which each agent has a state $(a)$, where $a$ is an integer and the discrepancy between states is $\delta$. A protocol using only averaging arrives at a configuration with a discrepancy between states of at most $2$ after time $O(\log\delta+\log n)$ whp.
Theorem 6.
Our majority protocol detects an anomaly with negligible probability. In the absence of an anomaly, the protocol’s output is always correct.
Proof.
Let bias $\Delta=|X|-|Y|$ and ${\cal S}_{r}$ be the sum of states $(a)$ in which each agent starts round $r$. Note that ${\cal S}_{1}=\Delta$. The averaging operation does not change the sum of agent states $(a)$ in round $r$. Thus as long as there are no $(x_{win})$ or $(y_{win})$ states, ${\cal S}_{r}$ is also the sum of states $(a)$ in which each agent ends round $r$. Due to doubling operation ${\cal S}_{r+1}=2{\cal S}_{r}$, as long as states $(x_{win})$ and $(y_{win})$ are not present.
We consider two cases. In the first case, one of the sets $X$ or $Y$ is larger than the other. Without loss of generality, assume this set is $X$. If there are no $(x_{win})$ or $(y_{win})$ states in round $r$, then ${\cal S}_{r}=2^{r-1}\Delta\geq 2^{r-1}$ and also ${\cal S}_{r}\leq 2n$. Note that if no $(x_{win})$ or $(y_{win})$ states had appeared until round $LOG+3$, then the sum ${\cal S}_{LOG+3}$ would be at least $2^{LOG+2}$ which is greater than $2n$. This proves that $(x_{win})$ or $(y_{win})$ states appear at the latest in round $LOG+3$.
We will now describe the dominant scenario that occurs whp. At the end of each round $r$, before the state $(x_{win})$ appears, the discrepancy between the values $(a)$ is at most $2$ whp due to Theorem 5. Furthermore, since ${\cal S}_{r}>0,$ some values $(a)$ must be positive. If eventually all values are (a)=(1|0|1)(a)=(-1|0|1), then they will double during transfer to round $r+1$. If eventually all values are $(a)=(0|1|2)$ and some value $(a)=(2)$, then the state $(x_{win})$ is created in round $r+1$. As we have already mentioned, the state $(x_{win})$ must appear in round $r+1\leq LOG+3$. In round $r+1$, the state $x_{win}$ is propagated to all agents in the population through the epidemic. This causes all agents in round $LOG+4$ to adopt state $(X_{win})$.
Now, let’s assume that no clock or protocol anomaly was detected. As mentioned before, in some round $r\leq LOG+3$, state $(x_{win})$ or $(y_{win})$ must have occurred. So the conclusion $|X|=|Y|$ cannot be reached. Since no anomaly was detected, at the end of round $r$, the leader must have also adopted the same state $(x_{win})$ or $(y_{win})$, as well as every other agent. Therefore, at the beginning of round $r$, all agents must have had state $(x_{win})$ or state $(0|1|2)$. Since $\Delta>0$, all states in any round cannot be equal to $(y_{win})$ or (2|1|0)(-2|-1|0). Thus, the outcome $|X|<|Y|$ cannot be reached. This guarantees that all agents adopt the correct final state $(X_{win})$ meaning $|X|>|Y|$.
In the second case, bias $\Delta=0$. In the dominant scenario, which holds whp, at the end of each round $r<LOG+4$, there are three possible states (a)=(1|0|1)(a)=(-1|0|1) due to Theorem 5. When moving to round $LOG+4$, the leader adopts the value $(tie)$, and in turn all other agents.
Now, let’s assume that no anomaly was detected. Assume also by contradiction that in some round $r<LOG+4$, state $(x_{win})$ or $(y_{win})$ appeared – without loss of generality, let this be $(x_{win})$. Then, since $\Delta=0$, state $(y_{win})$ or (2|1)(-2|-1) must also have appeared in round $r$. However, this would result in the detection of an anomaly in round $r+1$. This proves that state $(x_{win})$ or $(y_{win})$ did not appear before round $LOG+4$. This means that when moving to round $LOG+4$, both the leader and all other agents will assume the state $(tie)$, which reflects the correct outcome $|X|=|Y|$. ∎
3.5 System of Weights
In this section, we formulate the MC+ protocol, which creates weights that are disjoint sets of agents. Weight $W_{i}$ is a set consisting of $2^{i}$ agents. The weight creation rounds are indexed by integers $r=0,1,2,\ldots,LOG+2$. In round $r$, the protocol creates $W_{r}$, consisting of $2^{r}$ agents. If each of these agents is assigned unit mass, then $W_{r}$, composed of $2^{r}$ agents, has a total mass of $2^{r}$. At the start of the protocol, all followers are in state $\phi$, which marks free agents. At the beginning of each subsequent round $r$, if the leader has not yet reached a final state $max(r)$ or $Max(r)$, it adopts state $\lambda_{r}$, which serves as the precursor of $W_{r}$ carrying its mass $2^{r}$. The leader then transfers this mass to the first free agent it encounters via transition: $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(r)$, where $\lambda$ denotes a non-final leader state that carries no mass. Agents participating in the creation of $W_{r}$ have states $w_{r}(i)$, indicating that the agent possessing it belongs to $W_{r}$ and holds mass $2^{i}$ from that weight. Creating the weight is accomplished by a transition that divides the mass greater than $2^{0}$ between two agents: $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$. This process completes when all agents in $W_{r}$ have mass $1$, that is, when they reach state $w_{r}(0)$.
1:A set of agents with a single leader
2:Weights created among non-leaders
3:function WeightCreation($X$)
4:  The leader adopts state $\lambda$ and non-leaders adopt state $\phi$.
5:  for round $r\leftarrow 0$ to $LOG+2$ do
6:   Upon entering round $r$
7:      the leader changes state $\lambda$ to state $\lambda_{r}$ (with mass $2^{r}$).
8:      a follower changes local state $w_{r-2}(0)$ to global state $w_{r}$ and local $w$.
9:   The leader transfers mass $2^{r}$ to the first free agent it encounters:
10:      $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(r)$
11:   Distribute the mass by repeating: $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$ for $i>0$
12:   if state $w_{r-1}(i)$ for $i>0$ or $\lambda_{r-1}$ is present upon transition to round $r$ then
13:      $(w_{r-1}(i)\;|\;\lambda_{r-1})\rightarrow max(r-1)$
14:   Epidemic of state max(r1)max(r-1) to all agents which are in round $r$.
15:   Upon entering round $r$, leader changes state max(r2)max(r-2) to state Max(r2)Max(r-2).
16:   if a non-leader enters round $r$ interacting with an agent with state Max(r2)Max(r-2) then
17:      it adopts state Max(r2)Max(r-2).
18:   if a non-leader in state $w_{r-3}$ enters round $r$ interacting with state Max(r2)Max(r-2) then
19:      it marks state $w_{r-3}$ as top.
20:   Upon entering round $r$, agent changes state Max(r3)Max(r-3) to state endend.
21:   if a non-leader in state max(r2)max(r-2) enters round $r$
22:       interacting with state $\neq$ Max(r2)Max(r-2) then
23:      signal Anomaly
24:  Compare $W_{Max(r)-1}$ and $X$ using the Majority protocol
25:  (elements of $W_{Max(r)-1}$ have states $w_{Max(r)-1}$ marked as top)
26:  if $|W_{Max(r)-1}|<|X|$ then signal Anomaly
Algorithm 3 WeightCreation Protocol
Lemma 5.
Assume that at the beginning of round $r$ all $W_{i},$ for $i\leq r-1,$ have been created and the population has at least $2^{r+1}$ agents in state $\phi$. Then $W_{r}$ will be created in time $O(\log n)$ whp.
Proof.
There are exactly $2^{r}$ agents that do not have state $\phi$ at the start of round $r$. These agents include the leader and those involved in all weights created up to round r1r-1. The probability that an agent in state $w_{r}(i)$ encounters an agent in state $\phi$ during an interaction is at least $\frac{2}{n}\cdot\frac{2^{r+1}-2^{r}}{2^{r+1}+2^{r}}\geq\frac{2}{3n}$.
The process of creating $W_{r}$ can be modeled as a binary tree, with the root representing state $w_{r}(r)$. This state has two children, each of mass $2^{r-1}$, corresponding to state $w_{r}(r-1)$. At the next level, there are four states $w_{r}(r-2)$, each with mass $2^{r-2}$, and so on. At the final level, there are $2^{r}<n$ states $w_{r}(0)$, each of unit mass, represented as leaves.
Consider a path from the root to a leaf in this tree. This path consists of $r$ edges, each representing a transition $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$. We show that all transitions along any fixed path are completed in time $O(\log n)$ whp. The probability of executing the next transition on the path in a given interaction is at least $\frac{2}{3n}$. Over $cn\log n$ interactions, the expected number of interactions between the current agent on the path and an agent in state $\phi$ is at least $\frac{2}{3}c\log n$. By the Chernoff bound, this number exceeds $\log n>2^{r}$ whp for sufficiently large $c$. Thus, each transition along any fixed path completes in time $O(\log n)$ whp. By applying the union bound over all paths, we conclude that all transitions across all paths complete in time $O(\log n)$ whp. Consequently, the construction of $W_{r}$ also completes in time $O(\log n)$ whp. ∎
As there is a limit $n$ on the number of agents in the population, in some round $r$ there will not be enough agents to create the next $W_{r}$ in time $O(\log n)$. Our weight creation process will detect the first round $r_{max}$ in which the weight creation fails. It will be detected when an agent in state $\lambda_{r}$ or $w_{r}(i)$ for $i>0$ receives a signal to move to round $r+1$. In this case, the agent advancing to round $r+1$ will adopt state $max(r)$. This state is then propagated through the epidemic in round $r+1$ whp. The leader in state $max(r)$ when moving to round $r+2$ changes state to $Max(r)$. Then all other agents proceeding to round $r+2$ get state $Max(r)$. All these operations proceed as described above whp, calculating the value of $Max(r)$. The value of $Max(r)$ is the number assigned to the first weight that could not be created. If a follower reaches round $r+2$ with state $w_{r}(0)$, it is saved as an element of $W_{r}$ in the global part of the state. An anomaly is detected if an agent in state $max(r)$ interacts with an agent in round $r+2$ that is not in state $Max(r)$. The state $Max(r)$ occurs only in round $r+2$ and then transitions to the state endend. However, agents in the heaviest weight created are marked as top. This allows the population to know which weight is the highest, maintaining the desired number of states.
Theorem 7.
WeightCreation protocol (Algorithm 3) creates $W_{0},W_{1},W_{2},\ldots,W_{Max(r)-1}$ correctly with high probability, for some value $Max(r)>\log\frac{n}{3}$, also determining this value. With negligible probability, it correctly creates $W_{0},W_{1},W_{2},\ldots,W_{Max(r)-1}$, for some $Max(r)\leq\log\frac{n}{3},$ also determining this value. Otherwise, it detects an anomaly.
Proof.
By Lemma 5, the duration of the round is sufficient to produce $W_{r}$ whp when $2^{r}\leq n/3$. This ensures that all weights $r\leq\log\frac{n}{3}$ are created correctly whp. With negligible probability, the first incompletely formed $W_{r}$ satisfies $r\leq\log\frac{n}{3}$.
Let $W_{r}$ be the first weight that is not fully formed in its round. This occurs when, upon transitioning to round $r+1$, the state $\lambda_{r}$ or $w_{r}(i),$ for some $i>0,$ persists. Consequently, the state $max(r)$ emerges upon entering round $r+1$. This state is propagated whp to all agents in the population via an epidemic process during round $r+1$. When the leader advances to round $r+2$, it updates the state $max(r)$ to $Max(r)$. This causes all other agents to adopt state $Max(r)$ at the start of the round $r+2$ whp. With negligible probability, state $max(r)$ fails to reach the leader in round $r+1$. Assume also, no clock anomaly occurs. Then, the leader may advance to round $r+2$ without establishing the state $Max(r)$. This, in turn, triggers a protocol anomaly when an agent in state $max(r)$ progresses to round $r+2$. For pseudocode of Algorithm 3 see page 3. ∎
With negligible probability, the weight creation protocol may terminate prematurely. In such cases, the generated weights may be insufficient to execute the parity protocol for $X$. To ensure this does not occur in the current protocol execution, we use our majority protocol to compare the size of $X$ with the size of the largest created weight, Max(r)1Max(r)-1, after weight creation. The outcome $|X|<2^{Max(r)-1}$ guarantees that the remaining weights are sufficient to compute the parity for $X$.
3.6 Proof of Theorem 1
We start with the proof of Theorem 1, which states that there exists a population protocol for the parity problem that uses $O(\log^{3}n)$ states and silently stabilises in expected $O(\log^{3}n)$ time.
Proof of Theorem 1.
We can use the scheme described in Parity protocol (Algorithm 1) to combine the MC+ protocols for leader election, majority, and weight creation, obtaining an efficient Monte Carlo protocol for parity. By Theorem 3, this combination is proved to be an MC+ protocol for parity. In order to implement this combination, the global state component contains a bit that specifies whether the agent belongs to a weight that contributes to set $Y$. This bit is turned on or off by the result of the majority operation stored in its final local state. The number of states used by this process is $O(\log^{3}n)$. This can be shown as follows. The leader in each of the MC+ protocols needs a fixed number of states except for the clock component, which requires $O(\log^{3}n)$ states. Non-leaders need $O(\log^{2}n)$ states in the clock component. Those that belong to the weights use $O(\log n)$ states in the global component to determine the weight number and a constant number of states in the local component. Non-leaders that are not in the weights have a constant number of states in the global component and a logarithmic number in the local component.
By Theorem 2, the fast MC+ parity protocol can be combined with a slow parity protocol, resulting in an efficient, stable parity protocol. This protocol is silent because the slow parity protocol is silent. The total number of rounds of the MC+ parity protocol is $R=O(\log^{2}n)$. By Lemma 2, the slow parity protocol requires polynomial time. By Theorem 2, the expected running time of the fast stable parity protocol is $O(\log^{3}n)$. The total number of states in the combined parity protocol is equal to the sum of their numbers of states, i.e., $O(\log^{3}n)$.
4 The Congruence problem
In this section, we show how our solution to the parity problem naturally extends to deciding congruences modulo any integer $m$. The main challenge is constructing a polynomial-time silent protocol for the general congruence problem. We chose not to adapt the succinct congruence protocol from Czerner et al. [19], as it remains unclear whether this protocol is silent. Font shape ‘T1/lmr/m/scit’ undefinedOur fast congruence protocol follows the same structure as in the parity case.
4.1 Slow, Stable and Silent Congruence Protocol
In this subsection, we address the $m$-congruence problem, which involves determining whether the size of a distinguished subpopulation $X$ of agents is congruent to zero modulo a fixed integer $m>1$. We now outline a slow but always stabilising protocol for the congruence problem modulo any integer $m\geq 2$. We use the fact that using the binary representation of $m$, one can define a sequence $m_{0},m_{1},\ldots,m_{k}$ such that $m_{0}=1,m_{k}=m$ and $k=O(\log m)$. This sequence satisfies for each $i$ one of the relations $m_{i+1}=2m_{i}$ or $m_{i+1}=m_{i}+1$ and will be the sequence of possible agent masses. At the beginning of the protocol, all agents from the set $X$ have mass $m_{0}=1$, and all other agents have mass 0. If $m_{i+1}=2m_{i}$ and i<k1i<k-1, the interaction of two agents of mass $m_{i}$ leads to one of them acquiring mass $m_{i+1}$ and the other one acquiring mass 0. When i=k1i=k-1, such an interaction reduces the mass of both agents to zero. If, in turn, $m_{i+1}=m_{i}+1$, then the interaction of agents with masses $m_{i}$ and $1$ leads to a result analogous to the one described before. A limitation of this approach is that, during execution, no agents with mass $1$ may be available when needed for the second interaction type. Therefore, we introduce another type of interaction, in which $m_{i+1}=m_{i}+1$ and an agent with mass $m_{i}$ interacts with an agent with mass $m_{j}\neq 1,j\leq i$. In this case, the first agent acquires mass $m_{i+1}$, and the second mass $m_{j}-1$. If $m_{j}-1=m_{j-1}$, this interaction does not produce a mass outside the allowed set of agent masses. However, if $m_{j}=2m_{j-1}$, we must include mass $m_{j}-1$ in the allowed set, preserving a total of $O(\log m)$ distinct masses. To handle these extra masses, we introduce a rule: when an agent with mass $m_{j}-1$ interacts with a mass 0 agent, they transition to masses $m_{j-1}$ and $m_{j-1}-1$, respectively. The protocol described above reduces the total mass of agents from $|X|$ to mass $\mu\in[0,m)$, such that $\mu\equiv_{m}|X|$. To inform all mass 0 agents of the protocol’s outcome, any agent with non-zero mass interacting with a mass 0 agent sets the latter to state $f$, indicating that the congruence does not hold. Finally, in interactions that result in a mass reduction, agents with mass 0 receive a special state $T$. As long as they do not encounter agents with non-zero mass, they set the states of agents without mass to the state $t$, indicating that the congruence holds.
A formal description of the slow stable population congruence protocol SlowCongruence follows. This protocol uses $O(\log m)$ states and a crude analysis guarantees that it stabilises in expected time $O(n^{2})$. This protocol is used as a backup for the fast MC+ congruence protocol.
One can verify using binary encoding of $m$ that there exists a sequence $m_{0},\ldots,m_{k}$ such that:
  • $m_{0}=1$, $m_{k}=m$,
  • $k=O(\log m),$
  • for any $i\in[1,k],$ either $m_{i+1}=2m_{i}$ or $m_{i+1}=m_{i}+1$.
The protocol defines the following states, for which we provide an intuitive description:
  • $(0,T)$: a strong opinion that the congruence relation is satisfied,
  • $(0,t)$: a weak opinion that the congruence relation is satisfied,
  • $(0,f)$: a weak opinion that the congruence relation is not satisfied,
  • $(m_{i},F)$ for each $i\in[0,k]$: the state carrying the mass $m_{i}$ with a strong opinion that the congruence is not satisfied.
For each $i\leq k,$ s.t., $m_{i}\neq m_{i-1}+1$, we add the state $(m_{i}-1,F)$. In the initial configuration:
  • all elements of $X$ are in state $(1,F)$,
  • all other agents are in state $(0,t)$
Next, we specify the transitions:
[…]
Let the mass of an agent be defined as the value of the first coordinate of its state. A state is said to be carrying mass if its mass is greater than zero. Furthermore, the mass of a configuration is equal to the sum of the individual masses of all agents. The definitions of the initial configuration and the transition function lead to the following fact.
Fact 1.
  1. 1.
    The mass of the initial configuration is equal to $|X|$.
  2. 2.
    The total mass cannot increase.
  3. 3.
    In one transition, the mass can either remain unchanged or decrease by $m$.
  4. 4.
    The total mass cannot become negative.
We divide the states of the form $(x,F)$ into main states and auxiliary states. We say that a state $(m_{i},F)$ for $i\in[1,k]$ is a main state. Each state $(x,F)$ that is not main is called an auxiliary state. Thus auxiliary states have form $(m_{i}-1,F)$, where $m_{i}\neq m_{i-1}+1$.
Lemma 6 is instrumental in establishing the time efficiency of the SlowCongruence protocol.
Lemma 6.
If the total mass of the population is at least $m$, SlowCongruence protocol reduces this mass by $m$ in the expected time $O(mn)$.
Proof.
We call a configuration $i$-configuration when $i$ is the highest index for which there is an agent with the main state $(m_{i},F)$ or the auxiliary state $(m_{i+1}-1,F)$. We prove the following claim.
Claim 1.
The expected time to reach for the first time $j$-configuration, when we start from $i$-configuration for $i<j$ is at most $1.5(m_{j}-1)n$.
Proof of Claim.
We demonstrate our claim by induction on $j$. Consider the first appearance of a $j$-configuration with no main states $(m_{j},F)$. After time of at most $n/2$, an agent in auxiliary state $(m_{j+1}-1,F)$ adopts state $(m_{j},F)$ during interaction with an agent in state $(0,x)$. Two cases occur.
In the first case, where $m_{j+1}=m_{j}+1$, the state transition to $(m_{j+1},F)$ occurs as soon as an agent in state $({m_{j},F)}$ interacts with another agent in main state. If there is no agent in the main state, it is created after the first interaction of an auxiliary state and state $(0,x)$, and this will take on average time at most $n/2$. Subsequently, there will be at least one agent in the main state other than the one in state $(m_{j},F)$. On average time at most $n/2$, an interaction between these agents will occur, which results in the adoption of state $(m_{j+1},F)$ by some agent. Thus, the total time until $(j+1)$-configuration appears is at most
$1.5(m_{j}-1)n+n/2+n/2+n/2=1.5(m_{j+1}-1)n.$
In the second case, where $m_{j+1}=2m_{j}$, for as long as there is exactly one agent in state $(m_{j},F)$ in the $j$-configuration, this agent will not participate in any meaningful interactions. Therefore, we can consider a configuration of all agents, excluding a single agent in state $(m_{j},F)$. By inductive assumption, this configuration becomes a $j$-configuration after an average time of at most $1.5(m_{j}-1)n$. And when this happens, it takes an average time at most $n/2$, for a second agent to appear in state $(m_{j},F)$. Interaction of two agents in this state causes one of them to adopt state $(m_{j+1},F)$ after an average time of at most $n/2$. Therefore, in this case, the transition to the $(j+1)$-configuration takes an average time of at most
$1.5(m_{j}-1)n+1.5(m_{j}-1)n+n/2+n/2+n/2=1.5(m_{j+1}-1)n.$
The last step of the proof is estimating the time in which a mass reduction by transition (30),(40), or (70) occurs when we start from (k1)(k-1)-configuration. This can be handled analogously to the two previously considered cases, distinguished by whether $m_{k}=m_{k-1}+1$ or $m_{k}=2m_{k-1}$. This concludes the proof of Lemma 6, i.e., the mass reduction is obtained on average after time at most 1.5(m1)n1.5(m-1)n starting from an arbitrary configuration. ∎
There are two possible outcomes of the congruence computation. In the first, congruence holds, and the mass of the configuration is eventually reduced to zero. In the second, the total mass is reduced to an integer value $\mu\in(0,m)$, invalidating the congruence. The following lemma holds.
Lemma 7.
Assume that at some point the total mass $\mu\in(0,m)$. Then, all states carrying mass are distinct main states $(m_{i},F)$ and, except for the state with the smallest mass, all have masses $m_{i},$ where $m_{i+1}=2m_{i}$.
Proof.
Let $i$ be the greatest index such that $m_{i}\leq\mu$. By Claim 1, the state $(m_{i},F)$ is formed after average time at most $1.5(m_{i}-1)n+n/2$. Note that if $\mu=m_{i}$, the agent in state $(m_{i},F)$ will eventually accumulate the entire mass of $\mu$.
If, in turn, $m_{i}<\mu$, then $m_{i+1}=2m_{i}$ must hold. After formation of the state $(m_{i},F)$, the remaining mass of agents is $\mu^{\prime}=\mu-m_{i}<m_{i}$.
One can observe that the agents contributing to the mass $\mu^{\prime}$ will not have any meaningful interactions with the agent in state $(m_{i},F)$. Therefore, in subsequent interactions, they will form a set of agents with masses $(m_{j},F)$ as in the body of the lemma. This will take the total time
$\sum_{j}(1.5(m_{j}-1)n+n/2)=O(mn).$
Theorem 8.
The SlowCongruence protocol solves the $m$-congruence problem using $O(\log m)$ states and stabilises to a silent configuration in expected time $O(n^{2})$.
Proof.
The mass of the initial configuration is reduced to $\mu\in[0,m)$ through $\lfloor n/m\rfloor$ reductions, each requiring, by Lemma 6, expected time $O(mn)$. Hence, the total time of mass reduction is $O(mn\cdot\lfloor n/m\rfloor)=O(n^{2})$. Then the carrying mass agents reach their final configuration in time $O(mn)$, as established in Lemma 7. Finally, after time $O(n\log n)$, the agents with mass 0 reach their final states, which can be analysed similarly to the Coupon Collector’s Problem. In conclusion, the total stabilisation time required to reach a silent configuration is $O(n^{2})$. The number of states is $O(\log m)$, which follows directly from the definition of the protocol. ∎
4.2 Fast Congruence Protocol
The fast, silently stabilising congruence protocol can be constructed analogously to the fast parity protocol. The only remaining component of $m$-Congruence protocol (Algorithm 5) yet to be defined is the mechanism used to generate weights with masses $m,2m,4m,8m,\ldots$, referred to as $m$-WeightCreation (Algorithm 4). It differs from WeightCreation protocol only in a minor detail. To enable this new subprotocol, we consider the sequence of numbers $1=m_{0},m_{1},m_{2},\ldots,m_{k}=m$, which is the same as in the SlowCongruence protocol, and further extend this sequence with values $m_{k+i}=m2^{i}$, for $i>0$. And, for any integer $i>0$, we still get either $m_{i}=2m_{i-1}$ or $m_{i}=m_{i-1}+1$.
1:A set of agents with a single leader
2:$m$-Weights created among non-leaders
3:function $m$-WeightCreation($X$)
4:  The leader adopts state $\lambda$ and non-leaders adopt state $\phi$.
5:  for round $r\leftarrow 0$ to $LOG+2$ do
6:   Upon entering round $r$
7:      the leader changes state $\lambda$ to state $\lambda_{r}$ (with mass $m_{k+r}=m2^{r}$).
8:      a follower changes local state $w_{r-2}(0)$ to global state $w_{r}$ and local $w$
9:   The leader transfers mass $m_{k+r}$ to the first free agent it encounters:
10:      $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(k+r)$
11:   Distribute the mass by repeating for $i>0$:
12:   if $m_{i}=2m_{i-1}$ then $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$
13:   else $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(0)$
14:   if state $w_{r-1}(i)$ for $i>0$ or $\lambda_{r-1}$ is present upon transition to round $r$ then
15:      $(w_{r-1}(i)\;|\;\lambda_{r-1})\rightarrow max(r-1)$
16:   Epidemic of state max(r1)max(r-1) to all agents which are in round $r$.
17:   Upon entering round $r$, leader changes state max(r2)max(r-2) to state Max(r2)Max(r-2).
18:   if a non-leader enters round $r$ interacting with an agent with state Max(r2)Max(r-2) then
19:      it adopts state Max(r2)Max(r-2).
20:   if a non-leader in state $w_{r-3}$ enters round $r$ interacting with state Max(r2)Max(r-2) then
21:      it marks state $w_{r-3}$ as top.
22:   Upon entering round $r$, agent changes state Max(r3)Max(r-3) to state endend.
23:   if a non-leader in state max(r2)max(r-2) enters round $r$
24:       interacting with state $\neq$ Max(r2)Max(r-2) then
25:      signal Anomaly
26:  Compare $W_{Max(r)-1}$ and $X$ using the Majority protocol
27:  (elements of $W_{Max(r)-1}$ have states $w_{Max(r)-1}$ marked as top)
28:  if $|W_{Max(r)-1}|<|X|$ then signal Anomaly
Algorithm 4 m-WeightCreation Protocol
We conclude with a theorem whose proof closely parallels that of Theorem 1.
Theorem 9.
There exists a population protocol for the $m$-congruence problem that uses $O(\log^{3}n)$ states and silently stabilises in expected time $O(\log^{3}n)$.
1: $X$ a set of agents
2:true if $|X|$ is divisible by $m$, otherwise false
3:function $m$-Congruence($X$)
4:  LeaderElection()
5:  $m$-WeightCreation(X)
6:  $Y\leftarrow\emptyset$
7:  for $l\leftarrow LOG$ downto 0 do
8:   Add weight $m\cdot 2^{l}$ to $Y$
9:   Compare $X$ and $Y$ using the Majority protocol
10:   if $|X|<|Y|$ then Remove weight $m\cdot 2^{l}$ from $Y$
11:  if $|X|=|Y|$ then return true else return false
Algorithm 5 m-Congruence Protocol
5 Conclusion
In this work, we propose a new computing paradigm for population protocols, enabling the design of the first efficient, stable, and silent protocol for the parity and congruence problem. This new computing paradigm integrates several tools, including population weights, clocking mechanism, and efficient anomaly detection switching control to slow but always correct solutions. The new paradigm facilitates universal design of efficient multistage stable population protocols, enabling, e.g., efficient computation of more general congruencies. The power of the system of weights lies in the implicit conversion between unary and binary representations. In fact, due to this conversion, our parity protocol could also be amended to compute and report in a binary format (where, e.g., the consecutive bits of any integer value are stored in dedicated $O(\log n)$ agents) any subpopulation or even the whole population size, a problem posed in [21].
This raises a natural question: is binary representation of integers, both for input and output, better suited for designing efficient and stable population protocols? Although an efficient, stable multiplication protocol using binary representation is conceivable, achieving the same with unary representation is far more challenging, yet feasible (details omitted in this version). The primary obstacle is constructing a slow but stable protocol. Although recent years have seen remarkable progress in fast population protocols, advances in slow (polynomial time) yet stable protocols have been mainly limited to succinct protocols [19]. Our work indicates that further research is needed in this direction.

### v3

@@
 In this variant, in addition to efficient state utilisation, one is also concerned with the stabilisation time, determined by the number of interactions required to reach the final configuration, divided by the population size $n$.
-We say that a protocol is time-efficient if it stabilises in polylogarithmic (in $n$) time, and space-efficient if it utilises polylogarithmic state space.
+We call a protocol efficient if it is time‑efficient, stabilising in polylogarithmic time in $n$, and space‑efficient, using a polylogarithmic state space.
 We also say that a protocol is stable if it stabilises with the correct answer with probability 1, and silent if agents do not change their states after stabilisation.
-For nearly two decades, population protocols have been extensively studied, yielding efficient solutions for central problems in distributed computing, including leader election, and majority computation, a predicate type in Presburger Arithmetic closely tied to population protocols [5, 7].
-Surprisingly, no protocols have achieved both time- and space-efficiency for congruency predicates, including parity computation, which are complementary in this arithmetic framework.
-This gap highlights a significant challenge in the field [16, 19, 21, 24].
-To address this gap, we explore the parity problem, where agents are tasked with computing the parity of the given sub-population size.
-Then we extend the solution for parity to compute congruences modulo an arbitrary $m$.
-Previous research on efficient population protocols has focused on protocols that minimise both stabilisation time and state utilisation for specific problems.
-In contrast, this work slightly relaxes this expectation, permitting protocols to place less emphasis on full optimisation and more on universality, robustness, and probabilistic guarantees.
-This allows us to propose a novel computing paradigm that integrates population weights (or simply weights), a robust clocking mechanism, and efficient anomaly detection coupled with a switching mechanism (which ensures slow but always correct solutions).
+For nearly two decades, population protocols have been extensively investigated, producing efficient solutions for central problems in distributed computing, notably majority computation, one of the two primary predicates in Presburger arithmetic.
+Yet the complementary class of congruency predicates has remained substantially more elusive [5, 7].
+This long-standing gap constitutes a major open challenge in the area.
+Namely, efficient (non-stable) solutions exist via Monte-Carlo simulation of register machines [6], but all known stable protocols for these predicates are slow, either exponential time [16, 24] or polynomial [19].
+In this paper, we present a new methodology that enables the design of both efficient and stable protocols for computing congruency predicates.
+As a first step, we focus on the parity problem, in which agents must compute the parity of a designated sub-population size.
+We then generalise the obtained technique to compute congruence modulo an arbitrary integer $m$.
+Our congruency protocols achieve efficiency through a modular construction, with particular emphasis placed on universality, robustness, and strong probabilistic guarantees.
+Central to this approach is the MC+ computing paradigm, which enables integration of population weights (or simply weights), a robust clocking mechanism, and effective anomaly detection with a switching mechanism that guarantees fallback to a slow yet always-correct stable protocol.
 This paradigm facilitates universal design of efficient multistage stable population protocols.
@@
 We conclude by discussing the impact of implicit conversion between unary and binary representations enabled by the weight system, with applications to other problems, including the computation and representation of (sub-)population sizes, a problem posed in [21].
-All algorithmic solutions presented in this paper are stable, silent, and guarantee stabilisation time with high probability (whp) defined as $1-n^{-\eta},$ for a constant $\eta>0.$
+All algorithmic solutions presented in this paper are stable, silent, and guarantee stabilisation time in expectation and with high probability (whp) defined as $1-n^{-\eta},$ for a constant $\eta>0.$
 1 Introduction
@@
 When two agents engage in a direct interaction, their states change according to the predefined transition function, which forms an integral part of the population protocol.
-The weakest assumptions in population protocols concern the fixed (constant-size) operational state space of agents, with the population size $n$, whether exact or approximate, not hard-coded into either the states or the transition function.
 It is assumed that a protocol starts in the predefined initial configuration of agents’ states representing the input, and it stabilises in one of the final configurations of states representing the solution to the considered problem.
@@
 More recently, the focus has shifted to the parallel time, or simply the time, defined as the sequential time divided by the size $n$ of the whole population.
-The (parallel) time reflects on the parallelism of simultaneous independent interactions of agents utilised in efficient population protocols that stabilise in time $O(\mbox{poly}\log n)$.
-All algorithmic solutions presented in this paper are stable (always correct), silent (agents do not change their states after stabilisation), and guarantee stabilisation time with high probability (whp) defined as $1-n^{-\eta},$ for a constant $\eta>0.$
+The (parallel) time reflects the parallelism arising from simultaneous independent agent interactions used in efficient population protocols, which stabilise in time $O(\text{poly}\log\,n)$ while also using polylogarithmic state space.
 In this paper, the rules of a transition function are denoted as $A+B\rightarrow A^{\prime}+B^{\prime}$, meaning that if there is an interaction initiated by any agent in state $A$, with any agent in state $B$, then the former changes its state to $A^{\prime}$, while the latter changes its state to $B^{\prime}$.
-Some of the protocols presented in this work operate on compound states, which means that an agent state is a concatenation of some sets of substates $S_{1},\ldots,S_{k}$, and defining these sets defines the full state space as $S=S_{1}\times\ldots\times S_{k}$.
-Encoding a state of an agent can be seen as labeling it by a number represented in a binary form.
-Encoding compound states can be achieved by utilising specified positions in binary representation for each substate set.
+All algorithmic solutions presented in this paper are stable (always correct), silent (agents do not change their states after stabilisation), and guarantee stabilisation time with high probability (whp) defined as $1-n^{-\eta},$ for an arbitrary fixed constant $\eta>0.$
 1.1 Related Work and Results
-Population protocol studies primarily follow two main research avenues.
-The first focuses on understanding the behavior of large populations driven by simple interaction rules.
-The second treats population protocols as computational systems, investigating their computational power based on factors such as the type of scheduler and the memory available to individual agents.
-This study refers to the latter, focusing on designing efficient population protocols with robust probabilistic guarantees.
-The original population protocol model [5] assumes that the number of states and the transition function are constant, i.e., independent of the population size $n$.
-However, for fundamental problems such as leader election [23], majority computation [1], and the evaluation of various functions and predicates [9], it has been demonstrated that stable, fixed-state protocols operating in sublinear time are infeasible.
-This limitation has motivated the study of population protocols in which the number of states is permitted to grow with $n$, enabling such problems to be solved efficiently, i.e., in polylogarithmic time.
-Prior work on efficient population protocols has emphasized minimizing both stabilisation time and state usage for specific problems.
-In contrast, this study adopts a broader approach, prioritizing universality, robustness, and probabilistic guarantees over full optimisation.
-We introduce a novel computing paradigm that incorporates population weights (or simply weights), a robust clocking mechanism, and efficient anomaly detection paired with a switching mechanism to ensure slow but consistently correct solutions.
-This new paradigm supports the universal design of efficient multistage protocols, which are stable (always correct), silent (agents don’t alter their states after stabilisation), and ensure stabilisation in time O(polylog n)O($\text${polylog }n) with high probability.
+Population‑protocol research largely follows two directions: analysing the behaviour of large populations with predefined interaction rules; and studying computational power of population protocols under various schedulers and memory constraints.
+This work belongs to the latter, focusing on designing efficient protocols with robust probabilistic guarantees.
+The original population‑protocol model [5] assumes a fixed set of states an transitions, independent of the population size $n$.
+However, for fundamental tasks such as leader election [23], majority computation [1], two central tools in our computing paradigm, and the evaluation of various functions and predicates [9], it is known that stable, fixed‑state significantly faster than linear protocols are not feasible.
+This limitation has motivated models in which the number of states may grow with $n$, enabling efficient solutions.
+Much of the prior work on efficient stable population protocols has focused on specific problems.
+This study adopts a broader perspective, emphasising universality, robustness, and probabilistic performance guarantees.
+We propose the MC+ paradigm, which integrates population weights, a resilient clocking mechanism, and effective anomaly detection paired with a switching strategy that falls back to a slower but always correct protocol when necessary.
+This paradigm enables the universal construction of efficient multistage protocols that are stable (always correct), silent (agents cease changing state after stabilisation), and achieve stabilisation in time $O(\text{poly}\log\,n)$ whp.
 In the remainder of this section, we present an overview of both established and emerging concepts that underpin the new paradigm and its associated algorithmic solutions.
@@
 Epidemic Process
-A (one-way) epidemic process [6] is a simple yet effective protocol in which each agent can be in one of two possible states, denoted as 0 (susceptible) or $1$ (infected), with the transition function consisting of a single significant rule $(x)+(y)\rightarrow(x)+(\max\{x,y\})$.
+A (one-way) epidemic process [6] is a simple yet effective protocol in which each agent can be in one of two possible states, denoted as 0 (susceptible) or $1$ (infected), with the transition function consisting of a single meaningful rule $(x)+(y)\rightarrow(x)+(\max\{x,y\})$.
 Lemma 1 (Lemma 2 in [6]).
@@
 The epidemic process establishes a fundamental $\Theta(\log n)$-time communication bound in population protocols.
-It lies at the core of numerous algorithmic tools and solutions in population protocols, including phase clocks [2, 6, 8, 22, 26] and, in turn, fast leader election [3, 15, 2, 14, 8, 25, 13], majority computation [4, 1, 2, 15, 29, 30, 11, 10, 22], and others.
-A recent paper [27] presents a generalisation of the epidemic process to the $k$-contact epidemic model, in which an agent becomes infected only after interacting with $k$ distinct infected agents.
+It lies at the core of numerous algorithmic tools and solutions in population protocols, including phase clocks [2, 6, 8, 22, 26] and, in turn, efficient leader election [3, 15, 2, 14, 8, 25, 13], majority computation [4, 1, 2, 15, 29, 30, 11, 10, 22], and others.
+Recent work [27] introduces the $k$-contact epidemic model, requiring $k$ distinct infected contacts for infection.
 Efficient Leader Election
@@
 Later, Alistarh et al.
-[1] explored general trade-offs between state and time complexity, establishing a separation between slowly stabilising protocols using $o(\log\log n)$ states and rapidly stabilising ones requiring $O(\log n)$ states.
-Subsequent advances improved these bounds: Bilke et al.
+[1] explored general trade-offs between state and time complexity, establishing a separation between slowly stabilising protocols using $o(\log\log n)$ states and rapidly stabilising ones requiring $\Omega(\log\log n)$ states.
+Subsequent advances improved upper bounds: Bilke et al.
 [15] achieved $O(\log^{2}n)$ time whp with $O(\log^{2}n)$ states; Alistarh et al.
@@
 [13] achieved the optimal expected $O(\log n)$ time.
-In this paper, a direct use of leader election via pairwise duels appears in the slow parity protocol, see Section 2.1.
-In fast protocols, the elected leader initiates weight creation and coordinates rounds and process synchronisation.
-Our leader election mechanism builds on coin-flip reduction across successive rounds [6, 8, 13], with round numbering ensuring that at least one leader survives.
-In the rare event of multiple surviving leaders, anomaly detection resolves the conflict in expected $O(n)$ time, see Section 3.3.
-Majority Computation
+In our efficient protocols, the elected leader initiates weight creation and coordinates rounds and synchronisation.
+Our leader-election mechanism builds on coin-flip reduction across successive rounds [6, 8, 13], with round numbering ensuring at least one leader survives.
+In the rare case that multiple leaders remain, anomaly detection resolves the conflict in expected $O(n)$ time (Section 3.3).
+Majority Computation and Weights
 Let $A$ and $B$ be subpopulations of $n$ agents.
-The majority problem is to determine whether $|A|>|B|$, $|A|<|B|$, or $|A|=|B|$.
+The majority problem, a predicate from Presburger Arithmetic, is to determine whether $|A|>|B|$, $|A|<|B|$, or $|A|=|B|$.
 For efficient majority computation, Angluin et al.
@@
 Also, Kosowski and Uznański [28] achieved $O(1)$ states and $O(\mathrm{polylog}\,n)$ time whp.
-In this paper, we use exact majority computation as a conceptual balance scale, see Figure 2, enabling comparison of sub-population cardinalities through appropriately chosen weight combinations.
-Like other modern majority protocols, ours also uses cancelling (opposite biased agents $\to$ unbiased) and splitting (biased $\to$ two biased), doubling the bias per phase until $n$.
-As indicated earlier, our protocols prioritise universality, robustness, and probabilistic guarantees over full optimisation.
+In this paper, we use exact majority computation as a conceptual balance scale based on system of weights formed of disjoint subpopulations $W_{i}$ of agents, where $|W_{i}|=2^{i},$ for all $i,\dots,\log(\frac{n}{6}).$ This enables comparison of sub-population cardinalities through appropriately chosen weight combinations.
+Like other majority protocols, ours also uses cancelling (opposite biased agents $\to$ unbiased) and splitting (biased $\to$ two biased), doubling the bias per phase until $n$.
 For details, see Section 3.4.
-Figure 1: Balance scale. [^2] This is an AI-generated picture.
 Clocks and Synchronisation
@@
 In this paper, synchronisation is achieved via round counting by each agent, reinforced by leader-attributed fine-grained interaction counting and epidemic waves.
-This precise time-stamping system enables protocols with efficient anomaly detection through execution-time inconsistencies, essential for the MC+ protocols.
-MC+ Protocols
+This precise time-stamping system enables protocols with effective anomaly detection through execution-time inconsistencies, essential for the MC+ protocols.
+MC+ Paradigm
 Our new computing paradigm relies on Monte Carlo protocols enhanced with anomaly detection, termed MC+ protocols.
-These either silently stabilise to the correct answer or detect efficiently an anomaly, triggering a slow (polynomial-time) but always correct solution.
+These either silently stabilise to the correct answer or detect effectively an anomaly, triggering a slow (polynomial-time) but always correct solution.
 We distinguish between clock anomalies, caused by desynchronisation, and protocol anomalies, which arise from interactions between agents in conflicting states.
 We show that MC+ protocols are suitable building blocks for constructing more complex solutions.
+Because our MC+ protocols for leader election and majority must satisfy additional properties, their complexity does not match the fastest known results.
+Pursuing optimal bounds here would complicate the presentation and distract from the conceptual contributions.
 For detail see Section 3.
@@
 In the population protocol model, integer values are represented in unary format, where each value corresponds to the cardinality of a subpopulation of agents in a designated state.
-This representation makes addition straightforward, i.e., fast and stable.
+This representation makes addition straightforward, i.e., efficient and stable.
 While it is well understood that fixed-state Monte Carlo population protocols can simulate a limited variant of register machines (recognising problems in LOGSPACE) [6], the efficient and stable computation of basic operations such as subtraction (multiplication and division) remains significantly more challenging.
@@
 Our Contribution
-While this paper primarily focuses on the parity problem, our new computing paradigm also enables efficient computation of congruences, by providing a suitable framework which facilitates universal design of efficient multistage stable population protocols (Section 3.2).
-Specifically, the first efficient parity (Theorem 1) and congruence (Theorem 9) protocols are proved correct, both using $O(\log^{3}n)$ states and achieving stabilisation in $O(\log^{3}n)$ time (Section 4).
-Both protocols are stable (always correct), silent (agents do not change their states after stabilisation), and guarantee stabilisation time whp.
-We conclude (Section 5) by discussing the impact of implicit conversion between unary and binary representations enabled by the weight system, with applications to other problems, including computation and representation of (sub-)population sizes.
-To streamline the presentation, we assume that the subpopulations of agents representing input, temporary, and output values are each at most $\frac{n}{6}$.
-However, our solutions remain correct and efficient even for larger fractions of $n$.
-We also assume that the upper estimate $LOG$ on $\log n$ is provided with constant-factor accuracy and incorporated into the transition function.
-Finally, we note that a recent study [19] constructs succinctly encoded slow (polynomial time) population protocols for Presburger arithmetic, including congruences modulo any integer $m$.
-However, since [19] does not address the silence property, for completeness we include here a natural extension of our parity protocol that achieves silence.
+While this paper focuses primarily on the parity problem, our MC+ paradigm also enables efficient computation of congruences by supporting the universal design of efficient multistage stable population protocols (Section 3.2).We present the first stable and silent protocols for both parity (Theorem 1) and congruence (Theorem 9), each using $O(\log^{3}n)$ states and stabilising in $O(\log^{3}n)$ time both in expectation and with high probability (Section 4).
+In Section 5, we discuss the impact of implicit unary-to-binary conversion via the weight system and its applications to other problems, including (sub-)population size computation and representation.
+To streamline the presentation, we assume each subpopulation (input, temporary, output) is at most $\frac{n}{6}$; A recent work [19] constructs succinctly encoded, slow (polynomial-time) population protocols for Presburger arithmetic, including modular congruences.
+Since [19] does not address silence, we include also our generic polynomial congruence protocol.
 2 The Parity Problem
 Recall that in the parity problem, agents are tasked with computing the parity of a given sub-population size.
-In this work, we present a stable population protocol that solves the parity computation problem in polylogarithmic time.
-A common strategy for designing such protocols involves running a fast Monte Carlo algorithm that solves the problem with high probability, complemented by a slower, stable backup when needed.
-This approach has proven effective for problems such as majority computation and leader election, as discussed in the introduction.
-We adopt the same methodology in our construction.
-Notably, our protocol ensures both stabilisation and silence.
-We begin (Section 2.1) by presenting a simple protocol that computes the parity of a given subpopulation $X$ of agents, using a fixed number of states.
-It always stabilises in time $O(n\log n)$, both in expectation and with high probability.
-To our knowledge, it represents the current state-of-the-art for stable parity computation and serves as the backup in our overall protocol design.
+In this work, we present a stable population protocol that solves the parity computation problem in polylogarithmic time, see Theorem 1.
+We begin by presenting a fixed-state protocol that computes the parity of a given subpopulation $X$.
+This protocol serves as the fallback mechanism and stabilizes in $O(n\log n)$ time, both in expectation and with high probability.
 2.1 Slow Stable Parity Protocol
@@
 Initially, all agents in subpopulation $X$ start in $L_{1}$; all others start in $F_{0}$.
-On the interaction of two leaders, the rule $L_{x}+L_{y}\rightarrow L_{(x+y)\bmod 2}+F_{0}$ applies, where one remains a leader (with updated parity), and the other becomes a follower.
+In the interaction of two leaders, the rule $L_{x}+L_{y}\rightarrow L_{(x+y)\bmod 2}+F_{0}$ applies, where one remains a leader (with updated parity), and the other becomes a follower.
 Eventually, a single leader remains, encoding the parity of $|X|$ in its state index.
@@
 ∎
-2.2 Fast Parity Protocol
-In this section, we provide an outline of the first fast, stable, and silent parity protocol (Algorithm 1).
+2.2 Efficient Parity Protocol
+We present here an outline of the first efficient, stable, and silent parity protocol (Algorithm 1).
 1: $X$, a set of agents
@@
 Balance scale
-The balance scale idea stems from recomputation of powers of 2 (due to limited space), as proposed in [6].
-Equipped with polylogarithmic memory we can afford initial construction of subpopulations $W_{i}$ of cardinality $2^{i},$ for all $i=0,1,\dots,\lfloor\log\frac{n}{3}\rfloor$.
-These subpopulations serve as weights in balance scales (Figure 2) implemented via the fast majority protocol.
+The balance scale idea stems from recomputation of powers of 2, as proposed in [6].
+Equipped with polylogarithmic memory, we can afford the initial construction of subpopulations $W_{i}$ of cardinality $2^{i},$ for all $i=0,1,\dots,\lfloor\log\frac{n}{3}\rfloor$.
+These subpopulations serve as weights in balance scales implemented via the efficient majority protocol.
 Using subpopulation $W_{i}$, we can determine whether $|X|-|Y|\geq 2^{i}$ by computing exact majority (with tie detection) between $|X|$ and $|Y|+|W_{i}|$.
@@
 Theorem 1.
-There exists a population protocol which solves the parity problem utilising $O(\log^{3}n)$ states and stabilises silently in expected $O(\log^{3}n)$ time.
-The proof of Theorem 1 is provided in Section 3.6, near the end of the main part of this paper.
+There exists an efficient population protocol which solves the parity problem utilising $O(\log^{3}n)$ states and stabilises silently in expected $O(\log^{3}n)$ time.
 3 Efficient Protocols with Anomaly Detection
 In this section, we develop methods for constructing efficient population protocols that silently stabilise with certainty.
-Specifically, assuming the existence of two protocols: a slow, state-efficient protocol that stabilises silently in polynomial time with probability 1, and a fast Monte Carlo type protocol possibly composed of subprotocols, we show how to integrate these two to obtain a protocol that is both efficient in time and guaranteed to stabilise silently.
+Specifically, assuming the existence of two protocols: a slow, state-efficient protocol that stabilises silently in polynomial time with probability 1, and an efficient Monte Carlo type protocol possibly composed of subprotocols, we show how to integrate these two to obtain a protocol that is both efficient in time and guaranteed to stabilise silently.
 Monte Carlo protocols typically aim to compute the desired solution with high probability, but the correctness of the solution remains uncertain after execution.
-Existing mechanisms for detecting fast protocol failures and transitioning to a consistently stabilising protocol have been developed on an ad hoc basis.
-In contrast, we propose a novel universal anomaly-detection mechanism that effectively integrates multiple Monte Carlo subprotocols with a reliable, though slower, always-correct solution.
-Specifically, this mechanism identifies efficiently scenarios necessitating a shift to a slower, always-stabilising protocol and facilitates the creation of complex population protocols by sequentially combining simpler ones.
+Existing mechanisms for detecting efficient protocol failures and transitioning to a consistently stabilising protocol have been developed on an ad hoc basis.
+In contrast, we propose a novel systematic anomaly-detection mechanism that effectively integrates multiple Monte Carlo subprotocols with a reliable, though slower, always-correct solution.
+Specifically, this mechanism identifies efficiently scenarios necessitating a shift to a slower, always-stabilising protocol and facilitates the creation of complex efficient population protocols by sequentially combining simpler ones.
 Additionally, the constructed protocols are silent.
-Our efficient (polylogarithmic time and space) protocols converge to the correct solution with high probability or signal potential errors via an efficient anomaly-detection mechanism.
-Such signal is treated as a failure of the fast protocol, even if it does not always indicate an actual error.
-Our universal strategy is to first run a fast Monte Carlo protocol that either stabilises with high probability or, upon detecting an anomaly, switches to a slower, always-stabilising fallback.
+Our efficient protocols, requiring only $O(\text{poly}\log\,n)$ time and space, converge to the correct output with high probability or reliably signal potential errors through a effective anomaly-detection mechanism that enables fallback to guaranteed correctness.
+Such signal is treated as a failure of the efficient protocol, even if it does not always indicate an actual error.
+Our universal strategy is to first run an efficient Monte Carlo protocol that either stabilises with high probability or, upon detecting an anomaly, switches to a slower, always-stabilising fallback.
 3.1 Clock Protocol
-Central to our design is the clock protocol, which acts as a clocking device and a command counter, tracking algorithmic progress and synchronizing all agents to the same point in time.
+Central to our design is the clock protocol, which acts as a clocking device and a command counter, tracking algorithmic progress and synchronising all agents to the same point in time.
 In rare cases of desynchronisation, it signals an anomaly.
@@
 To measure a round’s duration, a leader-maintained interaction counter tracks $c\log n$ interactions, triggering the transition to the next round.
+As in the rest of this paper, we do not focus on the best possible optimisation of the clock protocol, but on using the simplest possible mechanisms to simplify presentation.
 Components of the clock protocol
@@
 3.2 MC+ Protocols
+In this subsection, we define MC+ protocols, which compute the correct result with high probability.
+However, unlike regular MC protocols, they provide guarantees of correct task execution.
+They complete the computation after polylogarithmic time.
+If, for some reason, the result is incorrect, they signal an anomaly.
+An anomaly can also be signaled with a correct result, but its overall probability is negligible.
+Therefore, the absence of an anomaly guarantees the correct execution of the computation.
+After completing one MC+ protocol, it is possible to continue the computation using the next protocol, assuming the previous one has executed correctly.
+However, if the previous protocol performs the computation incorrectly, signaling an anomaly ensures a transition to a stable, but possibly slow backup.
+For example, the design of the leader election protocol prevents the disappearance of all leaders, which would cause the clock to stop, but it is possible that several leaders remain after its completion.
+The calculations of the subsequent protocols performed assuming the calculation of a single leader will not produce a sensible result, but the inevitable interaction of different leaders will lead to anomaly signaling switching the calculations to a backup.
+State space.
 This part demonstrates how to effectively integrate multiple population protocols, necessitating a more sophisticated state space structure.
@@
 Elements $A_{0},B_{0}$ remain unchanged.
-In this work, we focus on fast population protocols, specifically Monte Carlo protocols with anomaly detection (MC+ protocols).
+In this work, we focus on efficient population protocols, specifically Monte Carlo protocols with anomaly detection (MC+ protocols).
 These protocols either silently stabilise on the correct computational output whp or trigger an anomaly detection signal.
@@
 Theorem 2.
-If a fast MC+ protocol and a polynomial-time stable protocol exist for the same problem, they can be combined to yield a protocol that is both fast (in expected running time) and stable (with high probability).
+If an efficient MC+ protocol and a polynomial-time stable protocol exist for the same problem, they can be combined to yield a protocol that is both time-efficient and stable whp.
 Moreover, if the polynomial-time stable protocol is silent, the resulting combined protocol is also silent.
 Proof.
-Assume that the fast protocol completes its computations in $R$ rounds.
+Assume that the efficient protocol completes its computations in $R$ rounds.
 If not successful, it reports a protocol anomaly in at most $n^{a}$ average time with negligible probability less than $n^{-a}$, for suitably chosen protocol parameters and $a>0$.
@@
   • the average time needed to handle protocol anomaly times its probability: $(n^{a}+n^{a})\cdot n^{-a}=2$.
-Therefore, the expected execution time of the combination of fast and slow protocol is at most $O(R\log n)+1+2=O(R\log n).$ Moreover, the combination of the protocols is stable and silent.
+Therefore, the expected execution time of the combination of efficient and slow protocol is at most $O(R\log n)+1+2=O(R\log n).$ Moreover, the combination of the protocols is stable and silent.
 Theorem 3.
@@
 ∎
-3.3 Leader Election
-The leader election protocol proposed in this work builds upon a scheme that has been employed in several prior studies [6, 8, 13].
+3.3 MC+ Leader Election
+The MC+ protocol for leader election proposed in this work builds upon a scheme that has been employed in several prior studies [6, 8, 13].
 In this approach, the leaders perform coin flips in successive rounds.
@@
 Unlike previous protocols, our method associates the coin-flip outcome (heads or tails) with the round number, thereby ensuring that at least one leader always persists.
-Apart from clock anomalies that can be detected during the execution of the protocol, the only other potential failure arises when multiple leaders remain.
-This is identified as a protocol anomaly when two such leaders interact after the leader election has concluded.
-This detection occurs in the expected time $O(n)$.
+Although the protocol is less efficient than several leader election protocols known from the literature, this construction reflects our emphasis on clarity and conceptual simplicity.
+Adapting more sophisticated existing protocols to the MC+ framework would likely entail additional technical difficulties.
+Furthemore, leader election does not constitute the bottleneck for the efficiency of protocols for parity or congruence.
+Besides detectable clock anomalies, the only other potential failure is persistent multiple leaders, which is detected as a protocol anomaly upon interaction after leader election concludes.
+This detection occurs in expected time $O(n)$.
 1: $X$, a set of agents
@@
 ∎
-3.4 Exact Majority
-We now construct and analyse our majority computation protocol.
-This protocol operates over $LOG+4$ rounds and achieves a time complexity of $O(\log^{2}n)$.
+3.4 MC+ Exact Majority
+In this section, we construct and analyse our majority computation protocol.
+We hypothesise that the best known protocol for majority, achieving time $O(\log n)$, could be modified into the MC+ protocol, which runs in a single or fixed number of rounds.
+Using such a protocol for parity computation would reduce its overall time to $O(\log^{2}n)$.
+Due to the complexity of this task, we have abandoned its implementation in this work and leave it as an open problem.
+Our MC+ majority protocol operates over $LOG+4$ rounds and achieves a time complexity of $O(\log^{2}n)$.
 Throughout this construction, we assume that a unique leader has already been selected.
@@
 Theorem 5 (Theorem 1 in [12]).
-Suppose we are given an initial configuration in which each agent has a state $(a)$, where $a$ is an integer and the discrepancy between states is $\delta$.
-A protocol using only averaging arrives at a configuration with a discrepancy between states of at most $2$ after time $O(\log\delta+\log n)$ whp.
+Given an initial configuration where each agent has an integer state $(a)$ and the state discrepancy is $\delta$, any protocol using only averaging reaches a configuration with discrepancy at most $2$ after $O(\log\delta+\log n)$ time whp.
+The following theorem holds.
 Theorem 6.
@@
 Proof.
-Let bias $\Delta=|X|-|Y|$ and ${\cal S}_{r}$ be the sum of states $(a)$ in which each agent starts round $r$.
-Note that ${\cal S}_{1}=\Delta$.
+Let bias $\Delta=|X|-|Y|$ and ${\mathcal{S}}_{r}$ be the sum of states $(a)$ in which each agent starts round $r$.
+Note that ${\mathcal{S}}_{1}=\Delta$.
 The averaging operation does not change the sum of agent states $(a)$ in round $r$.
-Thus as long as there are no $(x_{win})$ or $(y_{win})$ states, ${\cal S}_{r}$ is also the sum of states $(a)$ in which each agent ends round $r$.
-Due to doubling operation ${\cal S}_{r+1}=2{\cal S}_{r}$, as long as states $(x_{win})$ and $(y_{win})$ are not present.
+Thus as long as there are no $(x_{win})$ or $(y_{win})$ states, ${\mathcal{S}}_{r}$ is also the sum of states $(a)$ in which each agent ends round $r$.
+Due to doubling operation ${\mathcal{S}}_{r+1}=2{\mathcal{S}}_{r}$, as long as states $(x_{win})$ and $(y_{win})$ are not present.
 We consider two cases.
@@
 ∎
-3.5 System of Weights
-In this section, we formulate the MC+ protocol, which creates weights that are disjoint sets of agents.
+3.5 MC+ Weight Creation
+In this section, we introduce the MC+ protocol, which creates weights that are disjoint sets of agents.
 Weight $W_{i}$ is a set consisting of $2^{i}$ agents.
@@
 At the start of the protocol, all followers are in state $\phi$, which marks free agents.
-At the beginning of each subsequent round $r$, if the leader has not yet reached a final state $max(r)$ or $Max(r)$, it adopts state $\lambda_{r}$, which serves as the precursor of $W_{r}$ carrying its mass $2^{r}$.
+At the beginning of each subsequent round $r$ the leader posessing state $\lambda$ adopts state $\lambda_{r}$, which serves as the precursor of weight $W_{r}$ carrying its mass $2^{r}$.
 The leader then transfers this mass to the first free agent it encounters via transition: $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(r)$, where $\lambda$ denotes a non-final leader state that carries no mass.
 Agents participating in the creation of $W_{r}$ have states $w_{r}(i)$, indicating that the agent possessing it belongs to $W_{r}$ and holds mass $2^{i}$ from that weight.
-Creating the weight is accomplished by a transition that divides the mass greater than $2^{0}$ between two agents: $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$.
-This process completes when all agents in $W_{r}$ have mass $1$, that is, when they reach state $w_{r}(0)$.
+Creating the weight is accomplished by a transition that divides the mass greater than $2^{0}=1$ between two agents: $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$ when $i>0$.
+This process completes when all agents in $W_{r}$ have state $w_{r}(0)$ carrying mass 1.
+In some round $r$, the weight $W_{r}$ will inevitably fail to unfold.
+This will lead to a transition to round $r+1$ with state $\lambda_{r}$ or $w_{r}(i)$ for $i>0$ present in the population.
+In the protocol, we mark such an event with a transition to state $m(1)$.
+We will propagate state $m(1)$ to the entire population via an epidemic within round $r+1$ whp.
+The leader moving to a new round $r+2$ in state $m(1)$ changes state to $m(2)$.
+Then, all followers moving to round $r+2$ in state $m(1)$ will receive state $m(2)$.
+We signal an anomaly if the leader does not receive state $m(1)$ during the epidemic in round $r+1$.
+This causes an anomaly to be signaled when a follower in state $m(1)$ attempts to transition to round $r+2$.
+Note that if state $m(1)$ does not appear when transitioning to round $r+1$, the weight $W_{r}$ can be stored in the global states.
+This is done when transitioning to round $r+2$.
+With negligible probability, too few weights may be created to perform the parity protocol.
+To detect such an event, after the weights are created, their total mass is compared to the number of elements in the set $X$ for which the parity protocol is being performed.
+If an insufficient number of weights are created, an anomaly is signaled.
+Note that the absence of weights with masses between that of the largest created weight and $LOG$, i.e., these agent sets are empty – does not hinder the general scheme of the parity protocol.
 1:A set of agents with a single leader
@@
 3:function WeightCreation($X$)
-4:  The leader adopts state $\lambda$ and non-leaders adopt state $\phi$.
-5:  for round $r\leftarrow 0$ to $LOG+2$ do
-6:   Upon entering round $r$
-7:      the leader changes state $\lambda$ to state $\lambda_{r}$ (with mass $2^{r}$).
-8:      a follower changes local state $w_{r-2}(0)$ to global state $w_{r}$ and local $w$.
-9:   The leader transfers mass $2^{r}$ to the first free agent it encounters:
-10:      $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(r)$
-11:   Distribute the mass by repeating: $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$ for $i>0$
-12:   if state $w_{r-1}(i)$ for $i>0$ or $\lambda_{r-1}$ is present upon transition to round $r$ then
-13:      $(w_{r-1}(i)\;|\;\lambda_{r-1})\rightarrow max(r-1)$
-14:   Epidemic of state max(r1)max(r-1) to all agents which are in round $r$.
-15:   Upon entering round $r$, leader changes state max(r2)max(r-2) to state Max(r2)Max(r-2).
-16:   if a non-leader enters round $r$ interacting with an agent with state Max(r2)Max(r-2) then
-17:      it adopts state Max(r2)Max(r-2).
-18:   if a non-leader in state $w_{r-3}$ enters round $r$ interacting with state Max(r2)Max(r-2) then
-19:      it marks state $w_{r-3}$ as top.
-20:   Upon entering round $r$, agent changes state Max(r3)Max(r-3) to state endend.
-21:   if a non-leader in state max(r2)max(r-2) enters round $r$
-22:       interacting with state $\neq$ Max(r2)Max(r-2) then
-23:      signal Anomaly
-24:  Compare $W_{Max(r)-1}$ and $X$ using the Majority protocol
-25:  (elements of $W_{Max(r)-1}$ have states $w_{Max(r)-1}$ marked as top)
-26:  if $|W_{Max(r)-1}|<|X|$ then signal Anomaly
+4:  for round $r\leftarrow 0$ to $LOG+2$ do
+5:   upon entering round $r$ interacting with an agent in state different than $m(2)$:
+6:      if $r=0$ then
+7:        the leader adopts state $\lambda_{0}$ and non-leaders adopt state $\phi$.
+8:      else
+9:        the leader changes state $\lambda$ to state $\lambda_{r}$ (with mass $2^{r}$),
+10:        a follower changes local state $w_{r-2}(0)$ to global state $w_{r-2}$ and local $w$,
+11:        the leader changes state $\lambda_{r-1}$ to state $m(1)$,
+12:        a follower changes state $w_{r-1}(i)$ to state $m(1)$, when $i>0$,
+13:        the leader changes state $m(1)$ to state $m(2)$,
+14:        a follower in state $m(1)$ signals Anomaly.
+15:   upon entering round $r$ interacting with an agent in state $m(2)$:
+16:      a follower adopts state $m(2)$.
+17:   if interacting agents are in round $r$ then
+18:       The leader transfers mass $2^{r}$ to the first free agent it encounters: $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(r)$,
+19:        Distribute the mass by repeating for $i>0$: $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$,
+20:        Epidemic of state $m(1)$: $m(1)+\text{\emph{any state}}\rightarrow m(1)+m(1)$.
+21:  Compare $\bigcup W_{k}$ and $X$ using the Majority protocol
+22:    (elements of $\bigcup W_{k}$ have global states $w_{k}$)
+23:  if $|\bigcup W_{k}|<|X|$ then signal Anomaly
 Algorithm 3 WeightCreation Protocol
@@
 ∎
-As there is a limit $n$ on the number of agents in the population, in some round $r$ there will not be enough agents to create the next $W_{r}$ in time $O(\log n)$.
-Our weight creation process will detect the first round $r_{max}$ in which the weight creation fails.
-It will be detected when an agent in state $\lambda_{r}$ or $w_{r}(i)$ for $i>0$ receives a signal to move to round $r+1$.
-In this case, the agent advancing to round $r+1$ will adopt state $max(r)$.
-This state is then propagated through the epidemic in round $r+1$ whp.
-The leader in state $max(r)$ when moving to round $r+2$ changes state to $Max(r)$.
-Then all other agents proceeding to round $r+2$ get state $Max(r)$.
-All these operations proceed as described above whp, calculating the value of $Max(r)$.
-The value of $Max(r)$ is the number assigned to the first weight that could not be created.
-If a follower reaches round $r+2$ with state $w_{r}(0)$, it is saved as an element of $W_{r}$ in the global part of the state.
-An anomaly is detected if an agent in state $max(r)$ interacts with an agent in round $r+2$ that is not in state $Max(r)$.
-The state $Max(r)$ occurs only in round $r+2$ and then transitions to the state endend.
-However, agents in the heaviest weight created are marked as top.
-This allows the population to know which weight is the highest, maintaining the desired number of states.
+The following theorem holds.
 Theorem 7.
-WeightCreation protocol (Algorithm 3) creates $W_{0},W_{1},W_{2},\ldots,W_{Max(r)-1}$ correctly with high probability, for some value $Max(r)>\log\frac{n}{3}$, also determining this value.
-With negligible probability, it correctly creates $W_{0},W_{1},W_{2},\ldots,W_{Max(r)-1}$, for some $Max(r)\leq\log\frac{n}{3},$ also determining this value.
+WeightCreation protocol (Algorithm 3) creates $W_{0},W_{1},W_{2},\ldots,W_{Max(r)}$ correctly with high probability, for some value $Max(r)>\log\frac{n}{6}$.
+With negligible probability, it correctly creates $W_{0},W_{1},W_{2},\ldots,W_{Max(r)}$, for some $Max(r)\leq\log\frac{n}{6}$.
 Otherwise, it detects an anomaly.
@@
 This occurs when, upon transitioning to round $r+1$, the state $\lambda_{r}$ or $w_{r}(i),$ for some $i>0,$ persists.
-Consequently, the state $max(r)$ emerges upon entering round $r+1$.
+Consequently, the state $m(1)$ emerges upon entering round $r+1$.
 This state is propagated whp to all agents in the population via an epidemic process during round $r+1$.
-When the leader advances to round $r+2$, it updates the state $max(r)$ to $Max(r)$.
-This causes all other agents to adopt state $Max(r)$ at the start of the round $r+2$ whp.
-With negligible probability, state $max(r)$ fails to reach the leader in round $r+1$.
+When the leader advances to round $r+2$, it updates the state $m(1)$ to $m(2)$.
+This causes all other agents to adopt state $m(2)$ at the start of the round $r+2$ whp.
+With negligible probability, state $m(1)$ fails to reach the leader in round $r+1$.
 Assume also, no clock anomaly occurs.
-Then, the leader may advance to round $r+2$ without establishing the state $Max(r)$.
-This, in turn, triggers a protocol anomaly when an agent in state $max(r)$ progresses to round $r+2$.
+Then, the leader may advance to round $r+2$ without establishing the state $m(2)$.
+This, in turn, triggers a protocol anomaly when a follower in state $(1)$ progresses to round $r+2$.
 For pseudocode of Algorithm 3 see page 3.
 ∎
-With negligible probability, the weight creation protocol may terminate prematurely.
-In such cases, the generated weights may be insufficient to execute the parity protocol for $X$.
-To ensure this does not occur in the current protocol execution, we use our majority protocol to compare the size of $X$ with the size of the largest created weight, Max(r)1Max(r)-1, after weight creation.
-The outcome $|X|<2^{Max(r)-1}$ guarantees that the remaining weights are sufficient to compute the parity for $X$.
-3.6 Proof of Theorem 1
-We start with the proof of Theorem 1, which states that there exists a population protocol for the parity problem that uses $O(\log^{3}n)$ states and silently stabilises in expected $O(\log^{3}n)$ time.
-Proof of Theorem 1.
+3.6 Efficient Parity Protocol – Proof of Theorem 1
+Proof.
 We can use the scheme described in Parity protocol (Algorithm 1) to combine the MC+ protocols for leader election, majority, and weight creation, obtaining an efficient Monte Carlo protocol for parity.
@@
 Non-leaders that are not in the weights have a constant number of states in the global component and a logarithmic number in the local component.
-By Theorem 2, the fast MC+ parity protocol can be combined with a slow parity protocol, resulting in an efficient, stable parity protocol.
+By Theorem 2, the efficient MC+ parity protocol can be combined with a slow parity protocol, resulting in an efficient, stable parity protocol.
 This protocol is silent because the slow parity protocol is silent.
@@
 By Lemma 2, the slow parity protocol requires polynomial time.
-By Theorem 2, the expected running time of the fast stable parity protocol is $O(\log^{3}n)$.
+By Theorem 2, the expected running time of the efficient stable parity protocol is $O(\log^{3}n)$.
 The total number of states in the combined parity protocol is equal to the sum of their numbers of states, i.e., $O(\log^{3}n)$.
 4 The Congruence Problem
-In this section, we show how our solution to the parity problem naturally extends to deciding congruences modulo any integer $m$.
-The main challenge is constructing a polynomial-time silent protocol for the general congruence problem.
-We chose not to adapt the succinct congruence protocol from Czerner et al.
-[19], as it remains unclear whether this protocol is silent.
-Font shape ‘T1/lmr/m/scit’ undefinedOur fast congruence protocol follows the same structure as in the parity case.
+In this section, we argue that our parity protocol naturally extends to deciding congruences modulo any integer $m$.
+Our efficient congruence protocol (Section 4.2) follows the same structure as in the parity case.
+We do not adapt the polynomial‑time succinct protocol of Czerner et al.
+[19], as its silence is unclear.
+Instead, in Section 4.1, we present a protocol that is both stable and silent.
 4.1 Slow, Stable and Silent Congruence Protocol
@@
 This protocol uses $O(\log m)$ states and a crude analysis guarantees that it stabilises in expected time $O(n^{2})$.
-This protocol is used as a backup for the fast MC+ congruence protocol.
+This protocol is used as a backup for the efficient MC+ congruence protocol.
 One can verify using binary encoding of $m$ that there exists a sequence $m_{0},\ldots,m_{k}$ such that:
-  • $m_{0}=1$, $m_{k}=m$,
-  • $k=O(\log m),$
+  • $m_{0}=1$, $m_{k}=m$, $k=O(\log m),$
   • for any $i\in[1,k],$ either $m_{i+1}=2m_{i}$ or $m_{i+1}=m_{i}+1$.
@@
 ∎
-4.2 Fast Congruence Protocol
-The fast, silently stabilising congruence protocol can be constructed analogously to the fast parity protocol.
+4.2 Efficient Congruence Protocol
+The efficient, silently stabilising congruence protocol can be constructed analogously to the efficient parity protocol.
 The only remaining component of $m$-Congruence protocol (Algorithm 5) yet to be defined is the mechanism used to generate weights with masses $m,2m,4m,8m,\ldots$, referred to as $m$-WeightCreation (Algorithm 4).
@@
 5:  for round $r\leftarrow 0$ to $LOG+2$ do
-6:   Upon entering round $r$
-7:      the leader changes state $\lambda$ to state $\lambda_{r}$ (with mass $m_{k+r}=m2^{r}$).
-8:      a follower changes local state $w_{r-2}(0)$ to global state $w_{r}$ and local $w$
-9:   The leader transfers mass $m_{k+r}$ to the first free agent it encounters:
-10:      $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(k+r)$
-11:   Distribute the mass by repeating for $i>0$:
-12:   if $m_{i}=2m_{i-1}$ then $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$
-13:   else $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(0)$
-14:   if state $w_{r-1}(i)$ for $i>0$ or $\lambda_{r-1}$ is present upon transition to round $r$ then
-15:      $(w_{r-1}(i)\;|\;\lambda_{r-1})\rightarrow max(r-1)$
-16:   Epidemic of state max(r1)max(r-1) to all agents which are in round $r$.
-17:   Upon entering round $r$, leader changes state max(r2)max(r-2) to state Max(r2)Max(r-2).
-18:   if a non-leader enters round $r$ interacting with an agent with state Max(r2)Max(r-2) then
-19:      it adopts state Max(r2)Max(r-2).
-20:   if a non-leader in state $w_{r-3}$ enters round $r$ interacting with state Max(r2)Max(r-2) then
-21:      it marks state $w_{r-3}$ as top.
-22:   Upon entering round $r$, agent changes state Max(r3)Max(r-3) to state endend.
-23:   if a non-leader in state max(r2)max(r-2) enters round $r$
-24:       interacting with state $\neq$ Max(r2)Max(r-2) then
-25:      signal Anomaly
-26:  Compare $W_{Max(r)-1}$ and $X$ using the Majority protocol
-27:  (elements of $W_{Max(r)-1}$ have states $w_{Max(r)-1}$ marked as top)
-28:  if $|W_{Max(r)-1}|<|X|$ then signal Anomaly
+6:   upon entering round $r$ interacting with an agent in state different than $m(2)$:
+7:      if $r=0$ then
+8:        the leader adopts state $\lambda_{0}$ and non-leaders adopt state $\phi$.
+9:      else
+10:        the leader changes state $\lambda$ to state $\lambda_{r}$ (with mass $m_{k+r}=m2^{r}$),
+11:        a follower changes local state $w_{r-2}(0)$ to global state $w_{r-2}$ and local $w$,
+12:        the leader changes state $\lambda_{r-1}$ to state $m(1)$,
+13:        a follower changes state $w_{r-1}(i)$ to state $m(1)$, when $i>0$,
+14:        the leader changes state $m(1)$ to state $m(2)$,
+15:        a follower in state $m(1)$ signals Anomaly.
+16:   upon entering round $r$ interacting with an agent in state $m(2)$:
+17:      a follower adopts state $m(2)$.
+18:   if interacting agents are in round $r$ then
+19:       The leader transfers mass $m_{k+r}$ to the first free agent it encounters:
+20:        $\lambda_{r}+\phi\rightarrow\lambda+w_{r}(k+r)$
+21:        Distribute the mass by repeating for $i>0$:
+22:        if $m_{i}=2m_{i-1}$ then $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(i-1)$
+23:        else $w_{r}(i)+\phi\rightarrow w_{r}(i-1)+w_{r}(0)$
+24:        Epidemic of state $m(1)$: $m(1)+\text{\emph{any state}}\rightarrow m(1)+m(1)$.
+25:  Compare $\bigcup W_{k}$ and $X$ using the Majority protocol
+26:  (elements of $\bigcup W_{k}$ have global states $w_{k}$)
+27:  if $|\bigcup W_{k}|<|X|$ then signal Anomaly
 Algorithm 4 m-WeightCreation Protocol
-We conclude with a theorem whose proof closely parallels that of Theorem 1.
-Theorem 9.
-There exists a population protocol for the $m$-congruence problem that uses $O(\log^{3}n)$ states and silently stabilises in expected time $O(\log^{3}n)$.
 1: $X$ a set of agents
@@
 Algorithm 5 m-Congruence Protocol
-5 Conclusion
-In this work, we propose a new computing paradigm for population protocols, enabling the design of the first efficient, stable, and silent protocol for the parity and congruence problem.
-This new computing paradigm integrates several tools, including population weights, clocking mechanism, and efficient anomaly detection switching control to slow but always correct solutions.
-The new paradigm facilitates universal design of efficient multistage stable population protocols, enabling, e.g., efficient computation of more general congruencies.
-The power of the system of weights lies in the implicit conversion between unary and binary representations.
-In fact, due to this conversion, our parity protocol could also be amended to compute and report in a binary format (where, e.g., the consecutive bits of any integer value are stored in dedicated $O(\log n)$ agents) any subpopulation or even the whole population size, a problem posed in [21].
-This raises a natural question: is binary representation of integers, both for input and output, better suited for designing efficient and stable population protocols?
-Although an efficient, stable multiplication protocol using binary representation is conceivable, achieving the same with unary representation is far more challenging, yet feasible (details omitted in this version).
-The primary obstacle is constructing a slow but stable protocol.
-Although recent years have seen remarkable progress in fast population protocols, advances in slow (polynomial time) yet stable protocols have been mainly limited to succinct protocols [19].
-Our work indicates that further research is needed in this direction.
+We conclude with a theorem whose proof closely parallels that of Theorem 1.
+Theorem 9.
+There exists a population protocol for the $m$-congruence problem that uses $O(\log^{3}n)$ states and silently stabilises in expected time $O(\log^{3}n)$.
+5 Conclusions
+In this work, we introduce the MC+ computing paradigm for population protocols, enabling, for the first time, efficient, stable, and silent protocols for parity and congruence.
+The paradigm combines three main components: population weights, a robust clocking mechanism, and effective anomaly detection supported by a fallback to a slow but always‑correct stable solution.
+This framework supports universally applicable multistage stable protocols, allowing, for instance, efficient computation of general congruences.
+The strength of the weight system lies in its implicit conversion between unary and binary representations.
+As a result, our parity protocol easily extends to compute and output (in binary) the size of any designated subpopulation, or even the entire population, with consecutive bits stored across $O(\log n)$ dedicated agents, thereby resolving a problem posed in [21].
+This naturally raises the question of whether binary input and output offer a fundamentally better basis for designing efficient, stable population protocols than unary representation.
+While an efficient, stable multiplication protocol in binary seems plausible, achieving the same in unary remains far more difficult.
+Moreover, despite substantial recent progress on efficient population protocols, advances on slow yet stable protocols have largely been limited to succinct designs [16, 24, 19].
+Our work indicates that this direction deserves much more attention.
+Finally, further refinements of the MC+ paradigm, especially more efficient stable solutions for leader election and majority computation, would be highly valuable.